### v1

# Distributed Constrained Online Nonconvex Optimization with Compressed Communication

Abstract
This paper considers distributed online nonconvex optimization with time-varying inequality constraints over a network of agents. For a time-varying graph, we propose a distributed online primal–dual algorithm with compressed communication to efficiently utilize communication resources. We show that the proposed algorithm establishes an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}})$ network regret bound and an $\mathcal{O}({T^{1-{\theta_{1}}/2}})$ network cumulative constraint violation bound, where $T$ is the number of iterations and ${\theta_{1}}\in({0,1})$ is a user-defined trade-off parameter. When Slater’s condition holds (i.e, there is a point that strictly satisfies the inequality constraints at all iterations), the network cumulative constraint violation bound is reduced to $\mathcal{O}({T^{1-{\theta_{1}}}})$. These bounds are comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication for distributed online convex optimization with (time-varying) inequality constraints. Finally, a simulation example is presented to validate the theoretical results.
I Introduction
Distributed online convex optimization offers a promising framework for modeling a variety of problems in dynamic, uncertain, and adversarial environments, with wide-ranging applications such as real-time routing in data networks and online advertisement placement in web search [1]. This framework can be understood as a structured repeated game with $T$ iterations between a network of agents and an adversary. Specifically, at each iteration $t$, each agent $i$ selects a decision ${x_{i,t}}\in\mathbb{X}$, where $\mathbb{X}\subseteq{\mathbb{R}^{p}}$ is a known convex set and $p$ is a positive integer. Upon selection, the local loss function ${f_{i,t}}$ is privately revealed to agent $i$ by the adversary. The goal of agents is to collaboratively minimize the network-wide accumulated loss, and the corresponding performance metric is network regret
$\frac{1}{n}\sum\limits_{i=1}^{n}{\Big{(}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}-\mathop{\min}\limits_{x\in\mathbb{X}}\sum\limits_{t=1}^{T}{{f_{t}}({{x}})}}\Big{)}},$
where ${f_{t}}(x):=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$ is the global loss function of the network at iteration $t$.
Various projection-based distributed online algorithms with sublinear regret have been proposed to solve the distributed online convex optimization problem, see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10], and recent survey paper [11]. For example, for the fix communication topology, the authors of [2] propose a projection-based distributed online subgradient descent algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. For strongly convex local loss functions, the authors of [3, 4] establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound. For time-varying communication topology, the authors of [5] propose a projection-based distributed online weighted dual averaging algorithm, and establish an $\mathcal{O}({\sqrt{T}})$ regret bound for general convex local loss functions. By utilizing proportional-integral distributed feedback on the disagreement among neighboring agents, the authors of [6] propose a projection-based distributed online proportional-integral subgradient descent algorithm, and establish an $\mathcal{O}\big{(}{\log(T)}\big{)}$ regret bound for strongly convex local loss functions.
The aforementioned distributed online algorithms rely on agents exchanging their local data with perfect communication. Consequently, these algorithms encounter significant limitations arising from communication bottlenecks, particularly in scenarios involving limited bandwidth and power resources. The limitations becomes more pronounced as the scale of the multi-agent network and the dimensionality of the exchanged data increase. To overcome the limitations, distributed online algorithms with compressed communication are studied for the fix communication topology in the literature, see [12, 13, 14], and recent survey paper [15]. For example, the authors of [12] propose a decentralized online gradient descent algorithm with compressed communication by introducing an auxiliary variable to estimate the neighbors’ decisions at each iteration. They establish an $\mathcal{O}(\sqrt{T})$ network regret bound for general convex local loss functions. Unlike the compression strategy employed in [12], the authors of [13, 14] introduce two auxiliary variables: the first serves the same purpose as the auxiliary variable in [12], while the second ensures that the first variable does not need to be exchanged. The compression strategy is effective when the communication topology is fixed. However, it becomes ineffective for a time-varying communication topology.
Note that inequality constraints are common in practical applications. However, performing projection operations onto such constraints can result in substantial computational and storage burdens. To deal with this challenge, the authors of [16] consider the idea of long term constraints proposed in [17], where inequality constraints are allowed to be violated temporarily, with the requirement that they are ultimately satisfied over the long term. This violation is measured by a performance metric named constraint violation where the projection onto the non-negative orthant is performed after summing the constraint functions over time. Accordingly, they propose a distributed online primal–dual algorithm and establish an $\mathcal{O}({{T^{1/2+c}}})$ regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ constraint violation bound for general convex local loss and constraint functions, where $c\in({0,1/2})$ is a user-defined parameter. The regret bound is further reduced to $\mathcal{O}({{T^{c}}})$ for strongly convex local loss functions. The authors of [18] use performance metric named cumulative constraint violation where the projection onto the non-negative orthant is performed before summing the constraint functions over time, which is proposed in [19]. Moreover, they establish an $\mathcal{O}\big{(}{{T^{\max\{{c,1-c}\}}}}\big{)}$ regret bound and an $\mathcal{O}({{T^{1-c/2}}})$ cumulative constraint violation bound with $c\in({0,1})$ for quadratic local loss functions and linear constraint functions. However, [16, 18] only consider static inequality constraints. The authors of [20] extend distributed online convex optimization with long-term constraints into the time-varying constraints setting. Moreover, the same network regret and cumulative constraint violation bounds as in [18] are established. However, the distributed online algorithms proposed in [18, 20] are unable to achieve reduced network cumulative constraint violation under Slater’s condition. Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [21], and can be leveraged to achieve reduced constraint violation, e.g., [22, 23]. Recently, the authors of [24] propose a novel distributed online primal–dual algorithm, and establish reduced network cumulative constraint violation bounds under Slater’s condition.
Unlike the aforementioned studies that focus on distributed online convex optimization, the authors of [25] investigate distributed online nonconvex optimization where local loss functions are nonconvex. To evaluate algorithm performance, they propose a novel regret metric based on the first-order optimality condition associated with the variational inequality. For this metric, the offline benchmark seeks a stationary point of the cumulative global loss functions across all iterations. Moreover, they establish an $\mathcal{O}(\sqrt{T})$ regret bound for general nonconvex local loss functions. However, [25] does not account for inequality constraints and uses perfect communication among agents.
Motivated by the above observations, this paper considers the distributed online nonconvex optimization problem with time-varying constraints. For a time-varying communication topology, we propose a distributed online primal–dual algorithm with compressed communication to efficiently utilize communication resources. Furthermore, base on several classes of appropriately chosen parameter sequences, we analyze how compressed communication influences network regret and cumulative constraint violation. The contributions are as follows.
  • $\bullet$
    To the best of our knowledge, this paper is among the first to consider (time-varying) inequality constraints for distributed online nonconvex optimization. Compared to [2, 3, 4, 5, 6, 7, 8, 9, 10, 12, 13, 14, 16, 18, 20, 24] which focus on distributed online convex optimization, we consider distributed online nonconvex optimization where the absence of the convexity assumption on local loss functions makes the analysis more challenging. Compared to [25] which investigates distributed online nonconvex optimization, we additionally consider time-varying inequality constraints, which complicate both algorithm design and performance analysis. Moreover, similar to [12, 13, 14], we use compressed communication instead of perfect communication used in [25]. Different from [12, 13, 14] which consider a fixed communication topology, we consider a time-varying communication topology.
  • $\bullet$
    For the scaling parameter sequence $\{s_{t}\}$ produced by $\{1/{t^{\theta_{2}}}\}$ with ${\theta_{2}}>{\theta_{1}}$ and ${\theta_{1}}\in({0,1})$, we show in Theorem 1 that the proposed algorithm establishes an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},1+{\theta_{1}}-{\theta_{2}}}\}}}})$ network regret bound under ${\theta_{1}}<{\theta_{2}}<1$ and an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}})$ network regret bound under ${\theta_{2}}\geq 1$, and establishes an $\mathcal{O}({T^{1-{\theta_{1}}/2}})$ network cumulative constraint violation bound. When ${\theta_{2}}\geq 1$, these bounds are the same as the results established in [20, 24] where the local loss functions are convex and perfect communication is used. When Slater’s condition holds, we further show in Theorem 1 that the proposed algorithm establishes an reduced $\mathcal{O}({T^{1-{\theta_{1}}}})$ network cumulative constraint violation bound. This bound is the same as the results established in [24].
  • $\bullet$
    For the scaling parameter sequence $\{s_{t}\}$ produced by $\{\mu^{t}\}$ with $\mu\in({0,1})$, we show in Theorem 2 that the proposed algorithm establishes an $\mathcal{O}(\sqrt{T})$ network regret bound and an $\mathcal{O}({T^{3/4}})$ network cumulative constraint violation bound. These bounds are the same as the results established in Theorem 1 when ${\theta_{1}}=1/2$ and ${\theta_{2}}\geq 1$. Moreover, the network regret bound is the same as the results established in [25] where compressed communication and inequality constraints are not considered, as well as the results established in [12, 13] where inequality constraints and nonconvex local loss functions are not considered. When Slater’s condition holds, we further show that in Theorem 2 that the proposed algorithm establishes an reduced $\mathcal{O}(\sqrt{T})$ network cumulative constraint violation bound. This bound is the same as the results established in Theorem 1 when ${\theta_{1}}=1/2$.
The remainder of this paper is organised as follows. Section II presents the problem formulation. Section III proposes the distributed online primal–dual algorithm with compressed communication, and analyze its network regret and cumulative constraint violation bounds without and with Slater’s condition, respectively. Section IV provides a simulation example to verify the theoretical results. Finally, Section V concludes this paper. All proofs can be found in Appendix.
Notations: All inequalities and equalities throughout this paper are understood componentwise. ${\mathbb{N}_{+}}$, $\mathbb{R}$, ${\mathbb{R}^{p}}$ and $\mathbb{R}_{+}^{p}$ denote the sets of all positive integers, real numbers, $p$-dimensional and nonnegative vectors, respectively. Given $m$ and $n\in{\mathbb{N}_{+}}$, $[m]$ denotes the set $\{{1,\cdot\cdot\cdot,m}\}$, and $[m,n]$ denotes the set $\{{m,\cdot\cdot\cdot,n}\}$ for $m<n$. Given vectors $x$ and $y$, ${x^{T}}$ denotes the transpose of the vector $x$, and $\langle{x,y}\rangle$ denotes the standard inner. ${\mathbf{0}_{p}}$ and ${\mathbf{1}_{p}}$ denote the $p$-dimensional column vector whose components are all 00 and $1$, respectively. $\mathrm{col}({q_{1}},\cdot\cdot\cdot,{q_{n}})$ denotes the concatenated column vector of ${q_{i}}\in{\mathbb{R}^{{m_{i}}}}$ for $i\in[n]$. ${\mathbb{B}^{p}}$ and ${\mathbb{S}^{p}}$ denote the unit ball and sphere centered around the origin in ${\mathbb{R}^{p}}$ under Euclidean norm, respectively. For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $[x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$. For a function $f$ and a vector $x$, $\nabla f(x)$ denotes the subgradient of $f$ at $x$.
II Problem Formulation
Consider the distributed online nonconvex optimization problem with time-varying constraints. At iteration $t$, a network of $n$ agents is modeled by a time-varying directed graph ${\mathcal{G}_{t}}=({\mathcal{V},{\mathcal{E}_{t}}})$ with the agent set $\mathcal{V}=[n]$ and the edge set ${\mathcal{E}_{t}}\subseteq\mathcal{V}\times\mathcal{V}$. $({j,i})\in{\mathcal{E}_{t}}$ indicates that agent $i$ can receive information from agent $j$. The sets of in- and out-neighbors of agent $i$ are $\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})=\{{j\in[n]|({j,i})\in{\mathcal{E}_{t}}}\}$ and $\mathcal{N}_{i}^{\text{out}}({{\mathcal{G}_{t}}})=\{{j\in[n]|({i,j})\in{\mathcal{E}_{t}}}\}$, respectively. Let $\{{{f_{i,t}}:\mathbb{X}\to\mathbb{R}}\}$ and $\{{{g_{i,t}}:\mathbb{X}\to{\mathbb{R}^{{m_{i}}}}}\}$ be the sequences of nonconvex local loss and convex local constraint functions, respectively, where ${m_{i}}$ is a positive integer and ${g_{i,t}}\leq{\mathbf{0}_{{m_{i}}}}$ is the local constraint. Each agent $i$ selects a local decisions $\{{{x_{i,t}}\in\mathbb{X}}\}$ without prior access to $\{{{f_{i,t}}}\}$ and $\{{{g_{i,t}}}\}$. Upon selection, the nonconvex local loss function $\{{{f_{i,t}}}\}$ and convex local constraint function $\{{{g_{i,t}}}\}$ are privately revealed to the agent. The goal of the agent is to choose the decision sequence $\{{{x_{i,t}}}\}$ for $i\in[n]$ and $t\in[T]$ such that both network regret
${\rm{Net}\mbox{-}\rm{Reg}}(T)$ $:=\frac{1}{n}\sum\limits_{i=1}^{n}\Big{(}\sum\limits_{t=1}^{T}{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}}\rangle}-\mathop{\inf}\limits_{x\in\mathcal{X}_{T}}\Big{\langle}{\sum\limits_{t=1}^{T}{\nabla{f_{t}}({{x_{i,t}}})},x}\Big{\rangle}\Big{)},$ (1)
and network cumulative constraint violation
${\rm{Net}\mbox{-}\rm{CCV}}(T)$ $:=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}},$ (2)
increase sublinearly, where ${f_{t}}(x)=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$ is the global loss function of the network at iteration $t$, ${\mathcal{X}_{T}}=\{{x:x\in\mathbb{X},{g_{t}}(x)\leq{\mathbf{0}_{m}},\forall t\in[T]}\}$ is the feasible set, and ${g_{t}}(x)={\rm{col}}\big{(}{{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)}\big{)}\in{\mathbb{R}^{m}}$ with $m=\sum\nolimits_{i=1}^{n}{{m_{i}}}$ is the global constraint function of the network at iteration $t$. Similar to existing literature on distributed online convex optimization with time-varying constraints, e.g., [20, 24], we assume that the feasible set ${\mathcal{X}_{T}}$ is nonempty for all $T\in{\mathbb{N}_{+}}$, ensuring the existence of the offline optimal static decision.
[25] proposes an individual regret metric ${\rm{Net}\mbox{-}\rm{Reg}}(T)=\mathop{\max}\limits_{x\in\mathbb{X}}\Big{(}{\sum\nolimits_{t=1}^{T}{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-x}\rangle}}\Big{)}$ for distributed online nonconvex optimization by utilizing the first-order optimality condition associated with the variational inequality. In this paper, we consider time-varying inequality constraints, which cause the feasible set to become ${\mathcal{X}_{T}}$ instead of $\mathbb{X}$. Furthermore, the objective is to optimize the network-wide accumulated loss over all iterations, rather than the local one as considered in [25]. Therefore, we made a slight modification to the form of the regret metric in [25], transforming it into the form presented in (1).
The following commonly used assumptions are made throughout this paper.
Assumption 1.
The set $\mathbb{X}$ is convex and closed. Moreover, it is bounded by $R(\mathbb{X})$, i.e., for any $x\in\mathbb{X}$
$\|x\|\leq R(\mathbb{X}).$ (3)
Assumption 2.
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, the subgradients $\nabla{f_{i,t}}(x)$ and $\nabla{g_{i,t}}(x)$ exist. Moreover, there exist constants ${G_{1}}$ and ${G_{2}}$ such that
$\|{\nabla{f_{i,t}}(x)}\|$ $\leq{G_{1}},$ (4a)
$\|{\nabla{g_{i,t}}(x)}\|$ $\leq{G_{2}},x\in\mathbb{X}.$ (4b)
Due to the convexity of the local constraint function ${{g_{i,t}}}$, from Assumption 2 and Lemma 2.6 in [26], for all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, we have
$\|{{g_{i,t}}(x)-{g_{i,t}}(y)}\|$ $\leq{G_{2}}\|{x-y}\|,x,y\in\mathbb{X}.$ (5)
Assumption 3.
For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant ${L}$ such that
$\|{\nabla{f_{i,t}}(x)-\nabla{f_{i,t}}(y)}\|\leq{L}\|{x-y}\|,x,y\in\mathbb{X}.$ (6)
Remark 1.
Note that we do not require the assumptions that the local loss function ${{f_{i,t}}}$ and local constraint function ${{g_{i,t}}}$ are uniformly bounded, whereas [20] imposes them, and [24] imposes the first one. In addition, Assumption 2 is commonly used for distributed online convex optimization with long-term constraints [16, 18, 20, 24]. Since the local constraint function ${{g_{i,t}}}$ is convex, Assumption 2 implies its Lipschitz continuity on $\mathbb{X}$. In contrast, the nonconvex nature of the local constraint function ${{f_{i,t}}}$ precludes the existence of such a Lipschitz continuity property. This fundamental distinction causes analytical challenges compared to distributed online convex optimization, particularly in establishing network regret bounds. Furthermore, Assumption 3 plays a crucial role in addressing the analytical challenges, which is also used in distributed online nonconvex optimization [25].
Different from the studies on distributed online convex optimization with compressed communication [27, 12, 13, 14] that consider a fix and undirected graph, we consider a time-varying directed graph. The following assumption on the graph is made, which is also used in [16, 20, 24].
Assumption 4.
For $t\in{\mathbb{N}_{+}}$, the time-varying directed graph $\mathcal{G}_{t}$ satisfies that
(i) There exists a constant $w\in({0,1})$ such that ${[{{W_{t}}}]_{ij}}\geq w$ if $({j,i})\in{\mathcal{E}_{t}}$ or $i=j$, and ${[{{W_{t}}}]_{ij}}=0$ otherwise.
(ii) The mixing matrix ${W_{t}}$ is doubly stochastic, i.e., ${\sum\nolimits_{i=1}^{n}{[{{W_{t}}}]}_{ij}}={\sum\nolimits_{j=1}^{n}{[{{W_{t}}}]}_{ij}}=1$, $\forall i,j\in[n]$.
(iii) There exists an integer $B>0$ such that the time-varying directed graph $({\mathcal{V},\cup_{l=0}^{B-1}{\mathcal{E}_{t+l}}})$ is strongly connected.
In this paper, we consider the scenario where the communication between agents is compressed by using a class of compressors with globally bounded absolute compression error, as given in the following.
Assumption 5.
The compressor $\mathcal{C}:{\mathbb{Y}}\to{\mathbb{Y}}$ with $\mathbb{Y}\in{\mathbb{R}^{p}}$ satisfies
${\mathbf{E}_{\mathcal{C}}}[{\|{\mathcal{C}(x)-x}\|_{d}^{2}}]\leq C,\forall x\in{\mathbb{Y}},$ (7)
for some real norm parameter $d\geq 1$ and constant $C\geq 0$. Here ${\mathbf{E}_{\mathcal{C}}}$ denotes the expectation over the internal randomness of the stochastic compression operator $\mathcal{C}$.
Assumption 5 covers the majority of popular compressors in machine learning and signal processing applications such as the deterministic quantization used in [28, 29] and the unbiased stochastic quantization used in [30, 31]. The same class of compressors satisfying Assumption 5 is also used in [32, 14, 33].
Next, Slater’s condition is present in the following.
Assumption 6.
There exists a point ${x_{s}}\in\mathbb{X}$ and a positive constant ${\varsigma_{s}}$ such that
${g_{t}}({{x_{s}}})\leq-{\varsigma_{s}}{\mathbf{1}_{m}},t\in{\mathbb{N}_{+}}.$ (8)
Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [21]. However, for nonconvex problems, Slater’s condition alone does not generally guarantee strong duality because the standard Lagrange dual often has a nonzero duality gap in nonconvex problems [34]. For distributed online convex optimization, [24] establishes reduced network cumulative constraint violation bounds under Slater’s condition. However, to the best of our knowledge, there are no studies to show that similar reductions for cumulative constraint violation bounds can be achieved in distributed online nonconvex optimization, even for the general constraint violation bounds.
III Distributed Online Primal–Dual Algorithm with
Compressed Communication
In this section, we propose a distributed online primal–dual algorithm with compressed communication by using the class of compressors satisfying Assumption 5, and analyze the performance of the algorithm without and with Slater’s condition, respectively.
III-A Algorithm Description
To achieve reduced network cumulative constraint violation, [24] proposes a distributed online primal–dual algorithm for distributed online convex optimization with time-varying constraints. Here, we first give a subgradient descent variant of this algorithm in the following:
${x_{i,t}}$ $=\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{z_{j,t}}},$ (9a)
${v_{i,t+1}}$ $={\gamma_{t}}{[{{g_{i,t}}({{x_{i,t}}})}]_{+}},$ (9b)
${z_{i,t+1}}$ $={\mathcal{P}_{\mathbb{X}}}({{x_{i,t}}-{\alpha_{t}}{{\omega}_{i,t+1}}}),$ (9c)
${{\omega}_{i,t+1}}$ $={\nabla}{f_{i,t}}({x_{i,t}})+\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},$ (9d)
where ${{\gamma_{t}}}$ is the regularization parameter; and ${{\alpha_{t}}}$ is the stepsize.
To implement the algorithm (9), at each iteration each agent $i$ need to receive the exact vector-valued variable $z_{j,t}$ from its neighbors. That results in a substantial amount of data exchange over all iterations, especially when the scale of the multi-agent network and the dimension $p$ are large. To improve communication efficiency, we use the compressor satisfying Assumption 5 to compress $z_{j,t}$. Note that directly using the compressed variable $\mathcal{C}({{z_{j,t}}})$ to replace $z_{j,t}$ in (9a) does not work due to the compression error. To reduce the compression error, an auxiliary variable $\hat{z}_{j,t}\in{\mathbb{R}^{p}}$ and a scaling parameter ${s_{t}}$ can be introduced. Instead of directly compressing $z_{j,t}$, one can compress the scaled difference ${({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}$, then multiply the compressed result by ${s_{t}}$ and add it back to ${{\hat{z}}_{j,t-1}}$ to replace $z_{j,t}$ in (9a). Thus, (9a) is modified as follows:
${x_{i,t}}=\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\hat{z}_{j,t}}},$ (10)
where
${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)}.$ (11)
In this way, the compression error is reduced. However, as stated in [32], at each iteration each agent $i$ needs to receive the exact vector-valued variable $\hat{z}_{j,t-1}$ from its neighbors due to the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$ in (10). Therefore, the improved algorithm does not enjoy the benefits of compression.
To deal with the dilemma, [32] introduces another auxiliary variable ${y_{j,t}}\in{\mathbb{R}^{p}}$ to remove the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$. Then, (9a) becomes the following form:
${x_{i,t}}={{\hat{z}}_{j,t}}-{y_{j,t}},$ (12)
where
${{\hat{z}}_{j,t}}$ $={{\hat{z}}_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)},$ (13a)
${y_{j,t}}$ $={y_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}-{s_{t}}\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}}.$ (13b)
Note that agent $i$ only requires the compressed data as presented in (13). The similar strategy is also used in [13, 14]. When the graph $\mathcal{G}_{t}$ is fixed, i.e., ${[{W_{t-1}}]_{ij}}={[{W_{t}}]_{ij}}$ for all $t\in[T]$, it is straightforward to check that ${y_{j,t}}={{\hat{z}}_{j,t}}-\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t}}}$ for $j\in[n]$ by mathematical induction. However, the critical result does not hold when the graph $\mathcal{G}_{t}$ is time-varying.
To handle this challenge, instead of introducing another auxiliary variable, we define $\hat{z}_{j,t}$ in (10) as a vector-valued variable stored in agent $i$ rather than its neighbors. The variable $\hat{z}_{j,t}$ is indexed by $j$ to maintain one-to-one correspondence with all agents. Consequently, its value no longer needs to be exchanged, and only the compressed data $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$ in (11) is required by agent $i$. In this way, the distributed online primal–dual algorithm with compressed communication is proposed, which is presented in pseudo-code as Algorithm 1.
Algorithm 1 Distributed Online Primal–Dual Algorithm with Compressed Communication
0:  non-increasing stepsize sequence $\{{\alpha_{t}}\}\subseteq({0,+\infty})$, non-increasing scaling parameter sequence $\{{{s_{t}}}\}\subseteq({0,+\infty})$, and non-decreasing regularization parameter sequence $\{{\gamma_{t}}\}\subseteq({0,+\infty})$.
0:  ${\hat{z}_{j,0}}={\mathbf{0}_{p}}$ for $j\in[n]$, and ${z_{i,1}}\in{\mathbb{X}}$, $\forall i\in[n]$.
  for $t=1,\cdot\cdot\cdot$ do
     for $i=1,\cdot\cdot\cdot,n$ in parallel do
        Broadcast $\mathcal{C}\big{(}{({{z_{i,t}}-{\hat{z}_{i,t-1}}})/{s_{t}}}\big{)}$ to $\mathcal{N}_{i}^{\text{out}}({{\mathcal{G}_{t}}})$ and receive $\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)}$ from $j\in\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})$.
        Update
${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)},j\in\big{\{}\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})\cup\{i\}\big{\}}.$ (14)
        Select
${x_{i,t}}=\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}{{\hat{z}}_{j,t}}}.$ (15)
        Observe $\nabla{f_{i,t}}({{x_{i,t}}})$, $\nabla{g_{i,t}}({{x_{i,t}}})$, and ${g_{i,t}}({{x_{i,t}}})$.
        Update
${v_{i,t+1}}$ $={\gamma_{t}}{[{{g_{i,t}}({{x_{i,t}}})}]_{+}},$ (16a)
${{\omega}_{i,t+1}}$ $=\nabla{f_{i,t}}({x_{i,t}})+\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},$ (16b)
${z_{i,t+1}}$ $={\mathcal{P}_{\mathbb{X}}}({{x_{i,t}}-{\alpha_{t}}{{\omega}_{i,t+1}}}).$ (16c)
     end for
  end for
  $\{x_{i,t}\}$.
It is worth noting that agent $j$ does not need to receive ${{\hat{z}}_{j,t-1}}$ in (14) from any other agents to compute $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$, as it naturally possesses this information. In Algorithm 1, we initialize ${\hat{z}_{j,0}}={\mathbf{0}_{p}}$ for $j\in[n]$, and thus ${\hat{z}_{j,1}}={s_{1}}{\cal C}({z_{j,1}}/{s_{1}})$, which is inherently known by agent $j$ since $z_{j,1}$ is precisely stored in agent $j$. More generally, at each iteration $t$, agent $j$ has access to both $z_{j,t}$ and $\hat{z}_{j,t-1}$, enabling it to independently determine $\hat{z}_{j,t}$ without requiring information from any other agents.
III-B Performance Analysis
In this section, we establish network regret and cumulative constraint violation bounds for Algorithm 1 in the following theorems without and with Slater’s condition, respectively. Firstly, we choose the scaling parameter sequence $\{s_{t}\}$ produced by $\{1/{t^{\theta_{2}}}\}$ in the following theorem.
Theorem 1.
Suppose Assumptions 1–5 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with
${\alpha_{t}}=\frac{{\alpha_{0}}}{{{t^{{\theta_{1}}}}}},{\gamma_{t}}=\frac{{{\gamma_{0}}}}{{{\alpha_{t}}}},{s_{t}}=\frac{{s_{0}}}{{{t^{{\theta_{2}}}}}},$ (17)
where ${\theta_{1}}\in({0,1})$, ${\alpha_{0}}>0$, ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$, ${s_{0}}>0$, and ${\theta_{2}}>{\theta_{1}}$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]=\left\{\begin{array}[]{l}\mathcal{O}({{T^{\max\{{1-{\theta_{1}},1+{\theta_{1}}-{\theta_{2}}}\}}}}),{\rm{if}}\;{\theta_{1}}<{\theta_{2}}<1,\\ \mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}}),{\rm{if}}\;{\theta_{2}}\geq 1,\end{array}\right.$ (20)
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]=\mathcal{O}({T^{1-{\theta_{1}}/2}}).$ (21)
Moreover, if Assumption 6 also holds, then
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-{\theta_{1}}}}).$ (22)
Due to the space limitations, we omit the proof, which can be found in the arXiv version
Remark 2.
We show in Theorem 1 that Algorithm 1 establishes sublinear network regret and cumulative constraint violation bounds as in (20)–(21). These bounds characterize the impact of compressed communication on the network regret and cumulative constraint violation bounds, which is captured by ${\theta_{2}}$. When ${\theta_{2}}\geq 1$, compressed communication does not affect either of the bounds, and they are the same as the state-of-the-art results established by the distributed online algorithms without compressed communication in [20, 24]. When ${\theta_{1}}<{\theta_{2}}<1$, compressed communication may enable the network regret bound to become larger due to $01-{\theta_{2}}>$ 0, but does not affect the network cumulative constraint violation bound due to ${\theta_{2}}>{\theta_{1}}$. In addition, when Slater’s condition holds, the network cumulative constraint violation bound is further reduced as in (22). The bound is unaffected by compressed communication and remains the same as the results established by the distributed online algorithm with perfect communication in [24].
We then choose the scaling parameter sequence $\{s_{t}\}$ produced by $\{\mu^{t}\}$, which is also adopted by the distributed algorithms in [32, 14].
Theorem 2.
Suppose Assumptions 1–5 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with
${\alpha_{t}}={\alpha_{0}}\sqrt{\frac{{{\Psi_{t}}}}{t}},{\gamma_{t}}=\frac{{{\gamma_{0}}}}{{{\alpha_{t}}}},{s_{t}}={s_{0}}{\mu^{t}},$ (23)
where ${\Psi_{t}}=\sum\nolimits_{k=1}^{t}{{\mu^{k}}}$, ${\alpha_{0}}>0$, ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$, ${s_{0}}>0$, and $\mu\in({0,1})$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]=\mathcal{O}(\sqrt{T}),$ (24)
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]=\mathcal{O}({T^{3/4}}).$ (25)
Moreover, if Assumption 6 also holds, then
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}(\sqrt{T}).$ (26)
Due to the space limitations, we omit the proof, which can be found in the arXiv version
Remark 3.
We show in Theorem 2 that Algorithm 1 establishes an $\mathcal{O}(\sqrt{T})$ network regret bound as in (24) and an $\mathcal{O}({T^{3/4}})$ cumulative constraint violation bound as in (25). These bounds are the same as the results established in (20)–(21) with ${\theta_{1}}=1/2$ and ${\theta_{2}}\geq 1$. Moreover, the network regret bound is the same as the results established in [25] where compressed communication and inequality constraints are not considered, and the results established in [12, 13] where inequality constraints and nonconvex local loss functions are not considered. In addition, when Slater’s condition holds, the network cumulative constraint violation bound is further reduced as in (26), which is the same as the results established in (22) with ${\theta_{1}}=1/2$.
IV Simulation Example
To evaluate the performance of Algorithm 1, we consider a distributed online localization problem with long-term constraints over a network of $100$ sensors as follows:
$\mathop{\min}\limits_{x}\;\;\;\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\frac{1}{4}}}{\big{|}{{{\|{{S_{i}}-x}\|}^{2}}-{D_{i,t}}}\big{|}^{2}},$ (27a)
$\;{\rm{s}}{\rm{.t}}{\rm{.}}\;\;\;\;x\in\mathbb{X},{\rm{}}{B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in[T].$ (27b)
In this problem, the sensors aim to cooperatively track a moving target. The position of sensor $i$ is denoted by $S_{i}\in{\mathbb{R}^{p}}$. Sensor $i$ measures the distance between the positions of the target and itself by ${D_{i,t}}={\|{{S_{i}}-{X_{0,t}}}\|^{2}}+{\tau_{i,t}}$ where ${X_{0,t}}\in{\mathbb{R}^{p}}$ and ${\tau_{i,t}}\in\mathbb{R}$ denote the position of the target and the measurement noise at iteration $t$, respectively. Moreover, each agent only has access to its distance measurement. In addition, the position of the target is required to remain within a designated safe region characterized by the set $\mathbb{X}$, and avoid prolonged deviations from the mission region characterized by time-varying linear inequality constraint ${B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}}$ with ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ and ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$. However, occasional deviations are permitted to accommodate obstacle avoidance and exploration requirements. Inspired by [35, 36, 37], the problem as in (27) is formulated to achieve the least squares estimator for the position of the target. The communication topology is modeled by a time-varying undirected graph. Specifically, at each iteration $t$, the graph is first randomly generated where the probability of any two sensors being connected is $\rho$. Then, to make sure that Assumption 4 is satisfied, we add edges $({i,i+1})$ for $i\in[24]$ when $t\in\{{4c+1}\}$, edges $({i,i+1})$ for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges $({i,i+1})$ for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges $({i,i+1})$ for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$. Moreover, let ${[{{W_{t}}}]_{ij}}=\frac{1}{n}$ if $({j,i})\in{\mathcal{E}_{t}}$ and ${[{{W_{t}}}]_{ii}}=1-\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}$.
In this paper, we show in Theorem 1 that, both without and with Slater’s condition, Algorithm 1 establishes the same network regret and cumulative constraint violation bounds as the state-of-the-art results on distributed online convex optimization with long-term constraints, established by the distributed online algorithms with perfect communication in [24]. To verify the theoretical results, we compare Algorithm 1 with the algorithm in [24]. We set $\rho=0.1$, $\mathbb{X}={[{-5,5}]^{p}}$, $p=2$, ${m_{i}}=2$, and randomly choose each component of ${{S_{i}}}$ from the uniform distribution in the interval $[{-10,10}]$. We assume that the position of the target evolves by
[…]
where ${{Q_{t}}}$ is randomly generated from Bernoulli distribution with a success probability of $0.5$, and ${X_{0,0}}={[{0.8,0.95}]^{T}}$. Moreover, ${\tau_{i,t}}$ is randomly generated from the uniform distribution from in the interval $[{0,0.001}]$. Furthermore, each component of ${B_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{0,2}]$, and each component of ${b_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{b,b+1}]$ with $b>0$. Note that $b>0$ guarantees Slater’s condition holds. Here we choose $b=0.01$. In addition, we select the following compressor for Algorithm 1.
Standard uniform quantizer:
$\mathcal{C}(x)=\Delta\Big{\lfloor}{\frac{x}{\Delta}+\frac{{{\mathbf{1}_{p}}}}{2}}\Big{\rfloor},$
where $\Delta$ is a positive integer. This compressor satisfies Assumption 5 with $d=\infty$ and $C={\Delta^{2}}/4$, which is also used in [32, 14, 29, 33]. Transmitting $\mathcal{C}(x)$ requires $pq$ bits if each integer is encoded using $q$ bits. Here we set $\Delta=1$ and $q=8$. The inputs of all these algorithms are listed in TABLE I.
TABLE I: Input of algorithms.
Algorithms Inputs
Algorithm 1 in this paper ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},{s_{t}}=1/t$
The algorithm in [24] ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},\psi(x)={\|x\|^{2}}/2$
Figure 1: Evolutions of network regret.
Figure 2: Evolutions of network cumulative constraint violation.
Figs. 1 and 2 illustrate the evolutions of network regret and cumulative constraint violation, respectively. As shown in Fig.1, our Algorithm 1 exhibits almost the same network regret as that of the algorithm in [24]. Similarly, Fig. 2 demonstrates that our Algorithm 1 also has almost the same network cumulative constraint violation as that of the algorithm in [24]. However, due to compressed communication, our Algorithm 1 requires significantly fewer bits than those required by the algorithm in [24]. These simulation results are consistent with the results in Theorem 1.
V Conclusions
This paper studied the distributed online nonconvex optimization problem with time-varying constraints. To better utilize communication resources, we proposed a distributed online primal–dual algorithm with compressed communication. More importantly, the algorithm was able to handle time-varying communication topologies. We showed that the algorithm established sublinear network regret and cumulative constraint violation bounds. Moreover, the network cumulative constraint violation bounds were further reduced when Slater’s condition held. These bounds were comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication, even in the context of distributed online convex optimization with (time-varying) inequality constraints. In the future, we plan to investigate distributed bandit nonconvex optimization with time-varying constraints since gradient information is unavailable in many real-world applications.
A. Useful Lemmas
We begin by presenting several preliminary results that will be utilized in the subsequent proofs.
Lemma 1.
(Equation (5.4.21) on page 333 in [38]) For any $x\in{\mathbb{R}^{p}}$, it holds that ${\|x\|_{d}}\leq\hat{p}\|x\|$ and $\|x\|\leq\tilde{p}{\|x\|_{d}}$, where $\hat{p}={p^{\frac{1}{d}-\frac{1}{2}}}$ and $\tilde{p}=1$ when $d\in[{1,2}]$, and $\hat{p}=1$ and $\tilde{p}={p^{\frac{1}{2}-\frac{1}{d}}}$ when $d>2$.
Lemma 2.
([39, 40]) Let ${W_{t}}$ denote the mixing matrix associated with a time-varying graph that satisfies Assumption 4. Then,
$\Big{|}{{{[{\Psi_{s}^{t}}]}_{ij}}-\frac{1}{n}}\Big{|}\leq\tau{\lambda^{t-s}},\forall i,j\in[n],\forall t\geq s\geq 1,$ (31)
where $\Psi_{s}^{t}={W_{t}}{W_{t-1}}\cdots{W_{s}}$, $\tau={({1-\omega/4{n^{2}}})^{-2}}>1$, and $\lambda={({1-\omega/4{n^{2}}})^{1/B}}\in({0,1})$.
Lemma 3.
([41]) Let $\mathbb{K}$ denote a nonempty closed convex subset of ${\mathbb{R}^{p}}$ and let $b$ and $c$ denote two vectors in ${\mathbb{R}^{p}}$. If ${x}={\mathcal{P}_{\mathbb{K}}}({b-c})$, then for all $y\in\mathbb{K}$,
$2\langle{x-y,c}\rangle\leq{\|{y-b}\|^{2}}-{\|{y-x}\|^{2}}-{\|{x-b}\|^{2}}.$ (32)
In addition, let $\Phi(y)={\|{b-y}\|^{2}}+2\langle{c,y}\rangle$, then we know $\Phi$ is a strongly convex function with convexity parameter $\sigma=2$ and $x=\mathop{\arg\min}\limits_{y\in\mathbb{K}}\Phi(y)$. Then,
$\|{x-b}\|\leq{\|c\|}$ (33)
holds.
Lemma 4.
If Assumption 4 holds. For all $i\in[n]$ and $t\in{\mathbb{N}_{+}}$, ${\hat{z}_{i,t}}$ generated by Algorithm 1 satisfy
$\|{{\hat{z}_{i,t}}-{{\bar{z}}_{t}}}\|\leq\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{\hat{z}_{j,1}}}\|}+\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}+\|{\varepsilon_{i,t-1}^{z}}\|+\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|},$ (34)
where ${{\bar{z}}_{t}}=\frac{1}{n}\sum\nolimits_{i=1}^{n}{{\hat{z}_{i,t}}}$, $\varepsilon_{i,t-1}^{z}={\hat{z}_{i,t}}-{x_{i,t-1}}$.
Proof.
From (15), we recall that
${x_{i,t}}=\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}{\hat{z}_{j,t}}}$ (35)
holds.
From (35) and $\varepsilon_{i,t-1}^{z}={\hat{z}_{i,t}}-{x_{i,t-1}}$, we have
${\hat{z}_{i,t}}=\sum\limits_{j=1}^{n}{{{[{{W_{t-1}}}]}_{ij}}{\hat{z}_{j,t-1}}}+\varepsilon_{i,t-1}^{z}.$ (36)
Then, by following the proof of Lemma 4 in [20], we conclude that (34) holds. ∎
Lemma 5.
Suppose Assumptions 1–2 and 4–5 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 and $y$ be an arbitrary point in $\mathbb{X}$, then
${\mathbf{E}_{\mathcal{C}}}[\frac{1}{n}\sum\limits_{i=1}^{n}{v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}+\frac{1}{n}\sum\limits_{i=1}^{n}{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]$ $\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\mathbf{E}_{\mathcal{C}}}[{v_{i,t+1}^{T}{g_{i,t}}(y)}]+\frac{1}{n}\sum\limits_{i=1}^{n}{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]$
$\;\;+\frac{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\Delta}}_{t}}}]}{n}+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X})}{s_{t+1}}}{{{\alpha_{t}}}},$ (37)
where
${\Delta_{i,t}}(y)=\frac{1}{{2{\alpha_{t}}}}({{{\|{y-{x_{i,t}}}\|}^{2}}-{{\|{y-{x_{i,t+1}}}\|}^{2}}}),$
${{\tilde{\Delta}}_{t}}=\sum\limits_{i=1}^{n}{({{G_{1}}+{G_{2}}\|{{v_{i,t+1}}}\|})}\|{{x_{i,t}}-{z_{i,t+1}}}\|-\sum\limits_{i=1}^{n}{\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{2{\alpha_{t}}}}}.$
Proof.
Since the local constraint function ${g_{i,t}}$ is convex, we have
${g_{i,t}}(y)\geq{g_{i,t}}(x)+\nabla{g_{i,t}}(x)({y-x}),\forall x,y\in\mathbb{X}.$ (38)
We have
${\mathbf{E}_{\mathcal{C}}}[\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle]$ $={\mathbf{E}_{\mathcal{C}}}[\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-{z_{i,t+1}}}\rangle]+{\mathbf{E}_{\mathcal{C}}}[\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-y}\rangle]$
$\leq{G_{1}}{\mathbf{E}_{\mathcal{C}}}[\|{{x_{i,t}}-{z_{i,t+1}}}\|]+{\mathbf{E}_{\mathcal{C}}}[\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-y}\rangle],$ (39)
where the inequality holds due to (4a).
For the second term on the right-hand side of (39), from (16b), we have
${\mathbf{E}_{\mathcal{C}}}[\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{z_{i,t+1}}-y}\rangle]$ $={\mathbf{E}_{\mathcal{C}}}\big{[}\big{\langle}{\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},y-{z_{i,t+1}}}\big{\rangle}\big{]}+{\mathbf{E}_{\mathcal{C}}}[\langle{{\omega_{i,t+1}},{z_{i,t+1}}-y}\rangle]$
$={\mathbf{E}_{\mathcal{C}}}\big{[}\big{\langle}{\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},y-{x_{i,t}}}\big{\rangle}\big{]}+{\mathbf{E}_{\mathcal{C}}}\big{[}\big{\langle}{\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},{x_{i,t}}-{z_{i,t+1}}}\big{\rangle}\big{]}$
$\;\;+{\mathbf{E}_{\mathcal{C}}}[\langle{{\omega_{i,t+1}},{z_{i,t+1}}-y}\rangle].$ (40)
Next, we find the upper bound of each term on the right-hand side of (40).
From ${v_{i,t}}\geq{\mathbf{0}_{{m_{i}}}}$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$ and (38), we have
${\mathbf{E}_{\mathcal{C}}}\big{[}\big{\langle}{\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},y-{x_{i,t}}}\big{\rangle}\big{]}\leq{\mathbf{E}_{\mathcal{C}}}[v_{i,t+1}^{T}{g_{i,t}}(y)]-{\mathbf{E}_{\mathcal{C}}}[v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})].$ (41)
From the Cauchy–Schwarz inequality and (4b), we have
${\mathbf{E}_{\mathcal{C}}}\big{[}\big{\langle}{\big{(}{\nabla}{{g_{i,t}}({x_{i,t}})}\big{)}^{T}{v_{i,t+1}},{x_{i,t}}-{z_{i,t+1}}}\big{\rangle}\big{]}\leq{G_{2}}{\mathbf{E}_{\mathcal{C}}}[\|{{v_{i,t+1}}}\|\|{{x_{i,t}}-{z_{i,t+1}}}\|].$ (42)
By applying (32) to the update (16c), we have
$\;\;\;\;\;{\mathbf{E}_{\mathcal{C}}}[\langle{{\omega_{i,t+1}},{z_{i,t+1}}-y}\rangle]$
$\leq\frac{1}{{2{\alpha_{t}}}}({{\mathbf{E}_{c}}[{{\|{y-{x_{i,t}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{z_{i,t+1}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}]})$
$=\frac{1}{{2{\alpha_{t}}}}({\mathbf{E}_{\mathcal{C}}}[{{\|{y-{x_{i,t}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{x_{i,t+1}}}\|}^{2}}]+{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{x_{i,t+1}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{z_{i,t+1}}}\|}^{2}}]$
$\;\;-{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}])$
$={\mathbf{E}_{\mathcal{C}}}[{\Delta_{i,t}}(y)]+\frac{1}{{2{\alpha_{t}}}}\Big{(}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{y-\sum\limits_{j=1}^{n}{{{[{{W_{t+1}}}]}_{ij}}{\hat{z}_{j,t+1}}}}\Big{\|}^{2}}\Big{]}-{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{z_{i,t+1}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}]}\Big{)}$
$\leq{\mathbf{E}_{\mathcal{C}}}[{\Delta_{i,t}}(y)]+\frac{1}{{2{\alpha_{t}}}}\Big{(}{\sum\limits_{j=1}^{n}{{{[{{W_{t+1}}}]}_{ij}}{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{\hat{z}_{j,t+1}}}\|}^{2}}]}-{\mathbf{E}_{\mathcal{C}}}[{{\|{y-{z_{i,t+1}}}\|}^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}]\Big{)},$ (43)
where the last equality holds due to (15); and the last inequality holds since ${W_{t+1}}$ is doubly stochastic and ${\|\cdot\|^{2}}$ is convex.
We have
${\mathbf{E}_{\mathcal{C}}}[{\|{y-{\hat{z}_{i,t+1}}}\|^{2}}]-{\mathbf{E}_{\mathcal{C}}}[{\|{y-{z_{i,t+1}}}\|^{2}}]$ $={\mathbf{E}_{\mathcal{C}}}[\langle{y-{{\hat{z}}_{i,t+1}}+y-{z_{i,t+1}},y-{{\hat{z}}_{i,t+1}}-y+{z_{i,t+1}}}\rangle]$
$\leq{\mathbf{E}_{\mathcal{C}}}[\|{y-{\hat{z}_{i,t+1}}+y-{z_{i,t+1}}}\|\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|]$
$\leq 4R(\mathbb{X}){\mathbf{E}_{\mathcal{C}}}[\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|],$ (44)
where the first inequality holds due to the Cauchy–Schwarz inequality; and the last inequality holds due to (3).
We have
${({{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|}]})^{2}}$ $\leq{{\tilde{p}}^{2}}{({{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|_{d}}]})^{2}}$
$\leq{{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{{\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|_{d}^{2}}}]$
$={{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t}}-{s_{t+1}}\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
$={{\tilde{p}}^{2}}s_{t+1}^{2}{\mathbf{E}_{\mathcal{C}}}[{\|{({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}-\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
$\leq{{\tilde{p}}^{2}}Cs_{t+1}^{2},$ (45)
where the first inequality holds due to Lemma 1; the second inequality holds due to the Jensen’s inequality; the first equality holds due to (14); and the last inequality holds due to Assumption 5.
Summing (39)–(45) over $i\in[n]$, dividing by $n$, using $\sum\nolimits_{i=1}^{n}{{{\left[{{W_{t}}}\right]}_{ij}}}=1$, $\forall t\in{\mathbb{N}_{+}}$, and rearranging terms yields (37). ∎
Lemma 6.
Suppose Assumptions 1–2 and 4–5 hold. For all $i\in[n]$, let $\{{{x_{i,t}}}\}$ be the sequences generated by Algorithm 1 with ${\gamma_{t}}={\gamma_{0}}/{\alpha_{t}}$, where ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$ is a constant. Then, for any $T\in{\mathbb{N}_{+}}$,
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4{\alpha_{t}}}}}\Big{]}}}\leq 2G_{1}^{2}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+2{\tilde{p}}\sqrt{C}R(\mathbb{X})\sum\limits_{t=1}^{T}{\frac{{{s_{t+1}}}}{{{\alpha_{t}}}}},\forall y\in{\mathcal{X}_{T}},$ (46a)
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{1}{2}}}{\mathbf{E}_{\mathcal{C}}}\Big{[}{{\frac{{v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}}{{{\gamma_{t}}}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{2{\gamma_{0}}}}}}\Big{]}\leq{\mathbf{E}_{\mathcal{C}}}[{{\Lambda_{T}}(y)}]+{{{\tilde{\Lambda}}_{T}}(y)},\forall y\in{\mathcal{X}_{T}},$ (46b)
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}\leq n{\vartheta_{1}}+{\tilde{\vartheta}_{2}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]},$ (46c)
$\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}]}}}\leq{{\tilde{\vartheta}}_{3}}+{{\tilde{\vartheta}}_{4}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}},$ (46d)
${\mathbf{E}_{\mathcal{C}}}[\|{{z_{i,t+1}}-{x_{i,t}}}\|]\leq{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}{\mathbf{E}_{\mathcal{C}}}[\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|]},$ (46e)
where
${\Lambda_{T}}(y)=\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{v_{i,t+1}^{T}{g_{i,t}}(y)}}{{{\gamma_{t}}}}}},$
${{\tilde{\Lambda}}_{T}}(y)=\sum\limits_{i=1}^{n}{\frac{{{{\|{y-{x_{i,1}}}\|}^{2}}}}{{2{\gamma_{0}}}}}+2n{G_{1}}R(\mathbb{X})\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}}+2n{\gamma_{0}}G_{1}^{2}\sum\limits_{t=1}^{T}{\frac{1}{{\gamma_{t}^{2}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{{s_{t}}},$
${\vartheta_{1}}=\frac{{2\tau}}{{\lambda({1-\lambda})}}\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|},{{\tilde{\vartheta}}_{2}}=\frac{{4-4\lambda+2n\tau}}{{1-\lambda}},$
${{\tilde{\vartheta}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\vartheta}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{\left({1-\lambda}\right)}^{2}}}}+32.$
Proof.
(i) Since ${g_{i,t}}(y)\leq{\mathbf{0}_{{m_{i}}}}$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$ when $\forall y\in{\mathcal{X}_{T}}$, summing (37) over $t\in[T]$ gives
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{-v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})+{\Delta_{i,t}}(y)+\frac{1}{n}{{\tilde{\Delta}}_{t}}}\Big{]}+2{\tilde{p}}\sqrt{C}R(\mathbb{X})\sum\limits_{t=1}^{T}{\frac{{{s_{t+1}}}}{{{\alpha_{t}}}}}}}.$ (47)
We have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{({{G_{1}}+{G_{2}}\|{{v_{i,t+1}}}\|})}}\|{{x_{i,t}}-{z_{i,t+1}}}\|$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\Big{(}{2G_{1}^{2}{\alpha_{t}}+2G_{2}^{2}{\alpha_{t}}{{\|{{v_{i,t+1}}}\|}^{2}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4{\alpha_{t}}}}}\Big{)}}}.$ (48)
From (16a), for all $\forall t\in{\mathbb{N}_{+}}$, we have
$\|{{v_{i,t+1}}}\|={\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|.$ (49)
From (49), (16a) and the fact that ${\varphi^{T}}{[\varphi]_{+}}={\|{{{[\varphi]}_{+}}}\|^{2}}$ for any vector $\varphi$, we have
$2G_{2}^{2}{\alpha_{t}}{\mathbf{E}_{\mathcal{C}}}[{\|{{v_{i,t+1}}}\|^{2}}]-{\mathbf{E}_{\mathcal{C}}}[v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})]$ $=2G_{2}^{2}{\alpha_{t}}\gamma_{t}^{2}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|^{2}}]-{\gamma_{t}}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|^{2}}]$
$=({2G_{2}^{2}{\gamma_{0}}-1}){\gamma_{t}}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|^{2}}]\leq 0,$ (50)
where the last equality holds due to ${\gamma_{t}}={\gamma_{0}}/{\alpha_{t}}$; and the inequality holds due to ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$.
Combining (47)–(48) and (50) yields (46a).
(ii) From the Cauchy–Schwarz inequality and (3), we have
$\langle{\nabla{f_{i,t}}({{x_{i,t}}}),y-{x_{i,t}}}\rangle\leq\|{\nabla{f_{i,t}}({{x_{i,t}}})}\|\|{y-{x_{i,t}}}\|\leq 2{G_{1}}R(\mathbb{X}),\forall y\in\mathbb{X}.$ (51)
Dividing (37) by ${\gamma_{t}}$, using (51), and summing over $t\in[T]$ yields
$\;\;\;\;\;\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}]}}{{{\gamma_{t}}}}}}$
$\leq{\mathbf{E}_{\mathcal{C}}}[{{\Lambda_{T}}(y)}]+{2n{G_{1}}R(\mathbb{X})}\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}}{{{\gamma_{t}}}}+}}\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\Delta}}_{t}}}]}}{{{\gamma_{t}}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X})}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{{s_{t+1}}}.$ (52)
We have
$\;\;\;\;\;\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{({{G_{1}}+{G_{2}}\|{{v_{i,t+1}}}\|})\|{{x_{i,t}}-{z_{i,t+1}}}\|}}{{{\gamma_{t}}}}}}$
$\leq\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{({\frac{{2{\gamma_{0}}G_{1}^{2}}}{{\gamma_{t}^{2}}}+\frac{{2{\gamma_{0}}G_{2}^{2}{{\|{{v_{i,t+1}}}\|}^{2}}}}{{\gamma_{t}^{2}}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4{\gamma_{0}}}}})}}.$ (53)
From ${\gamma_{t}}={\gamma_{0}}/{\alpha_{t}}$, we have
$\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}}{{{\gamma_{t}}}}}=\frac{1}{{2{\gamma_{0}}}}\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{y-{x_{i,t}}}\|}^{2}}-{{\|{y-{x_{i,t+1}}}\|}^{2}}}]}\leq\frac{{{{\|{y-{x_{i,1}}}\|}^{2}}}}{{2{\gamma_{0}}}}.$ (54)
From (49), (16a) and the fact that ${\varphi^{T}}{[\varphi]_{+}}={\|{{{[\varphi]}_{+}}}\|^{2}}$ for any vector $\varphi$, we have
$\frac{{2{\gamma_{0}}G_{2}^{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{v_{i,t+1}}}\|}^{2}}}]}}{{\gamma_{t}^{2}}}-\frac{{{\mathbf{E}_{\mathcal{C}}}[{v_{i,t+1}^{T}{g_{i,t}}({{x_{i,t}}})}]}}{{2{\gamma_{t}}}}$ $=2{\gamma_{0}}G_{2}^{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]-\frac{{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}{2}$
$=({2{\gamma_{0}}G_{2}^{2}-\frac{1}{2}}){\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]\leq 0,$ (55)
where the inequality holds due to ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$.
Combining (52)–(55), and noting that $\{{{s_{t}}}\}$ is nonincreasing, we have (46b).
(iii) From (15) and $\sum\nolimits_{i=1}^{n}{{{[{{W_{t}}}]}_{ij}}}=\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}=1$, we have
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{\sum\limits_{k=1}^{n}{{{[{{W_{t}}}]}_{ik}}{{\hat{z}}_{k,t}}}-{{\bar{z}}_{t}}+{{\bar{z}}_{t}}-\sum\limits_{k=1}^{n}{{{[{{W_{t}}}]}_{jk}}{{\hat{z}}_{k,t}}}}\Big{\|}}\Big{]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{\sum\limits_{k=1}^{n}{{{[{{W_{t}}}]}_{ik}}{{\hat{z}}_{k,t}}}-{{\bar{z}}_{t}}}\Big{\|}}\Big{]}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{{{\bar{z}}_{t}}-\sum\limits_{k=1}^{n}{{{[{{W_{t}}}]}_{jk}}{{\hat{z}}_{k,t}}}}\Big{\|}}\Big{]}}}$
$=2\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}{{\hat{z}}_{j,t}}}-{{\bar{z}}_{t}}}\Big{\|}}\Big{]}}=2\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{\|}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}({{{\hat{z}}_{j,t}}-{{\bar{z}}_{t}}})}}\Big{\|}}\Big{]}}$
$\leq 2\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}{\mathbf{E}_{\mathcal{C}}}[{\|{{{\hat{z}}_{j,t}}-{{\bar{z}}_{t}}}\|}]}=2\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{\hat{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}]}.$ (56)
We have
$\sum\limits_{t=3}^{T}{\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}}\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,s}^{z}}\|}]}=\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,t}^{z}}\|}]}\sum\limits_{s=0}^{T-t-2}{{\lambda^{s}}}}\leq\frac{1}{{1-\lambda}}\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,t}^{z}}\|}]}}.$ (57)
From (56), (34), (57), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\leq 2\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{\hat{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}]}}$
$\leq 2\tau\sum\limits_{t=1}^{T}{{\lambda^{t-2}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{\hat{z}_{j,1}}}\|}}}+\frac{2}{n}\sum\limits_{t=2}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,t-1}^{z}}\|}]}}}+2\sum\limits_{t=2}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t-1}^{z}}\|}]}}$
$\;\;+2\tau\sum\limits_{t=3}^{T}{\sum\limits_{i=1}^{n}{\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,s}^{z}}\|}]}}}}$
$\leq\frac{{2\tau}}{{\lambda({1-\lambda})}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{\hat{z}_{j,1}}}\|}}+4\sum\limits_{t=2}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t-1}^{z}}\|}]}}+\frac{{2n\tau}}{{1-\lambda}}\sum\limits_{t=1}^{T-2}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{j,t}^{z}}\|}]}}$
$=\frac{{2\tau}}{{\lambda({1-\lambda})}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\|{{\hat{z}_{j,1}}}\|}}+4\sum\limits_{t=1}^{T-1}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}+\frac{{2n\tau}}{{1-\lambda}}\sum\limits_{t=1}^{T-2}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}$
$\leq n{\vartheta_{1}}+{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}.$ (58)
Therefore, from (58), we know that (46c) holds.
(iv) Similar to the way to get (56), from (15) and $\sum\nolimits_{i=1}^{n}{{{[{{W_{t}}}]}_{ij}}}=\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}=1$, and ${\|\cdot\|^{2}}$ is convex, we have
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}]}}\leq 4\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{{{\hat{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}^{2}}]}.$ (59)
From (34), we have
     $\;\;\;\;\;4\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{{{\hat{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}^{2}}]}}$
$\leq 4{\sum\limits_{t=1}^{T}\sum\limits_{i=1}^{n}{\mathbf{E}_{\mathcal{C}}}\Big{[}\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{{\hat{z}}_{j,1}}}\|}+\|{\varepsilon_{i,t-1}^{z}}\|+\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}+\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}}\Big{)}^{2}}\Big{]}$
$\leq 16\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{{\hat{z}}_{j,1}}}\|}}\Big{)}^{2}}+{{\|{\varepsilon_{i,t-1}^{z}}\|}^{2}}+{\Big{(}{\frac{1}{n}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,t-1}^{z}}\|}}\Big{)}^{2}}+{\Big{(}{\tau\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}}\Big{)}^{2}}}\Big{)}}\Big{]}$
$\leq 16\sum\limits_{i=1}^{n}\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}\Big{[}\Big{(}{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{{\hat{z}}_{j,1}}}\|}}\Big{)}^{2}}+{{\|{\varepsilon_{i,t-1}^{z}}\|}^{2}}+\frac{1}{n}\sum\limits_{j=1}^{n}{{{\|{\varepsilon_{j,t-1}^{z}}\|}^{2}}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+{\tau^{2}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}{\Big{(}{\sum\limits_{j=1}^{n}{\|{\varepsilon_{j,s}^{z}}\|}}\Big{)}^{2}}}\Big{)}\Big{]}$
$\leq 16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{{\hat{z}}_{j,1}}}\|}}\Big{)}^{2}}+2{{\|{\varepsilon_{i,t-1}^{z}}\|}^{2}}+\frac{{n{\tau^{2}}}}{{1-\lambda}}\sum\limits_{s=1}^{t-2}{{\lambda^{t-s-2}}}\sum\limits_{j=1}^{n}{{{\|{\varepsilon_{j,s}^{z}}\|}^{2}}}}\Big{)}}}\Big{]}$
$=16\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}\Big{[}{\Big{(}{{\Big{(}{\tau{\lambda^{t-2}}\sum\limits_{j=1}^{n}{\|{{{\hat{z}}_{j,1}}}\|}}\Big{)}^{2}}+2{{\|{\varepsilon_{i,t-1}^{z}}\|}^{2}}}\Big{)}+\frac{{16{n^{2}}{\tau^{2}}}}{{1-\lambda}}\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T-2}{{{\|{\varepsilon_{j,t}^{z}}\|}^{2}}}}\sum\limits_{s=0}^{T-t-2}{{\lambda^{s}}}}}\Big{]}$
$\leq{{\tilde{\vartheta}}_{3}}+{{\tilde{\vartheta}}_{4}}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}]}},$ (60)
where the third inequality holds due to the Hölder’s inequality.
Therefore, from (59)–(60), we know that (46d) holds.
(v) Applying (33) to the update (16c) gives
${\mathbf{E}_{\mathcal{C}}}[\|{{z_{i,t+1}}-{x_{i,t}}}\|]\leq{{\alpha_{t}}{\mathbf{E}_{\mathcal{C}}}[\|{{\omega_{i,t+1}}}\|]}$ $={{\alpha_{t}}{\mathbf{E}_{\mathcal{C}}}[\|{\nabla{f_{i,t}}({{x_{i,t}}})+{{{\big{(}{\nabla{g_{i,t}}({{x_{i,t}}})}\big{)}}^{T}}}{v_{i,t+1}}}\|]}$
$\leq{{\alpha_{t}}}{\mathbf{E}_{\mathcal{C}}}[{\|{\nabla{f_{i,t}}({{x_{i,t}}})}\|+\|{{{{\big{(}{\nabla{g_{i,t}}({{x_{i,t}}})}\big{)}}^{T}}}{v_{i,t+1}}}\|}]$
$\leq{{\alpha_{t}}}{\mathbf{E}_{\mathcal{C}}}[{{G_{1}}+{G_{2}}{\gamma_{t}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]$
$={\mathbf{E}_{\mathcal{C}}}[{{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}],$ (61)
where the first equality holds due to (16b); the last inequality holds due to (4a), (4b), and (49); and last equality holds due to ${\gamma_{t}}={\gamma_{0}}/{\alpha_{t}}$.
Therefore, from (61), we know that (46e) holds. ∎
Lemma 7.
Under the same conditions as stated in Lemma 6, and supposing that Assumption 3 holds, for any $T\in{\mathbb{N}_{+}}$ and any $y\in{\mathcal{X}_{T}}$, it holds that
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}\leq 2nR(\mathbb{X})L{\vartheta_{1}}+{\vartheta_{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+2{\tilde{p}}\sqrt{C}R(\mathbb{X})\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}},$ (62a)
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}\leq\sqrt{{\vartheta_{3}}T+{\vartheta_{4}}T{{\tilde{\Lambda}}_{T}}(y)+4n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}T\sum\limits_{t=1}^{T}{s_{t}^{2}}},$ (62b)
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}\leq n{G_{2}}{\vartheta_{1}}+{\vartheta_{5}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+{\vartheta_{6}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+n{\tilde{p}}\sqrt{C}{G_{2}}{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}},$ (62c)
where
${\vartheta_{2}}=2G_{1}^{2}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2},{\vartheta_{3}}=2G_{2}^{2}{{\tilde{\vartheta}}_{3}},{\vartheta_{4}}=\frac{{4\max\{{1,2G_{2}^{2}{{\tilde{\vartheta}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},{\vartheta_{5}}={n{G_{1}}{G_{2}}{{\tilde{\vartheta}}_{2}}},$
${\vartheta_{6}}={1+G_{2}^{2}{\gamma_{0}}{{\tilde{\vartheta}}_{2}}}.$
Proof.
(i) From ${f_{t}}(x)=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$, we have
$\nabla{f_{t}}(x)=\frac{1}{n}\sum\nolimits_{j=1}^{n}{\nabla{f_{j,t}}(x)}.$ (63)
From (63), we have
$\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{t}}({{x_{i,t}}})}]}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{j,t}}({{x_{i,t}}})}]}}$
$=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{j,t}}({{x_{j,t}}})}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{j,t}}({{x_{i,t}}})-\nabla{f_{j,t}}({{x_{j,t}}})}]}}$
$=\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{i,t}}({{x_{i,t}}})}]}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{j,t}}({{x_{i,t}}})-\nabla{f_{j,t}}({{x_{j,t}}})}]}}.$ (64)
From (64) and (6), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{L\|{{x_{i,t}}-{x_{j,t}}}\|\|{{x_{i,t}}-y}\|}]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}},\forall y\in\mathbb{X},$ (65)
where last inequality holds due to (3).
For the second term on the right-hand side of (65), from (46c), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{\varepsilon_{i,t}^{z}}\|}]}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{z_{i,t+1}}-{{\hat{z}}_{i,t+1}}}\|}]}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{x_{i,t}}-{z_{i,t+1}}}\|}]}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{4nR{{(\mathbb{X})}^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}{\alpha_{t}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4n{\alpha_{t}}}}}\Big{]}}}$
$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}]}}{{4{\alpha_{t}}}}}},$ (66)
where the second inequality holds due to $\varepsilon_{i,t}^{z}={{\hat{z}}_{i,t+1}}-{x_{i,t}}$; and the third inequality holds since (45) holds and $\{{{s_{t}}}\}$ is nonincreasing.
Combining (65)–(66) and (46a), and noting that $\{{{s_{t}}}\}$ is nonincreasing, we know that (62a) holds.
(ii) We have
${\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]$ $={\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}+{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]$
$\geq\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]-{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]$
$\geq\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]-{\mathbf{E}_{\mathcal{C}}}[{{{\|{{g_{i,t}}({{x_{i,t}}})-{g_{i,t}}({{x_{j,t}}})}\|}^{2}}}]$
$\geq\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]-G_{2}^{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}},$ (67)
where the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
From ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$, we have
$\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]}}}=\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}.$ (68)
From (67)–(68), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}$
$\leq\frac{2}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}}+\frac{{2G_{2}^{2}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}}]}}}$
$\leq{\vartheta_{3}}+2\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}+2G_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{\varepsilon_{i,t}^{z}}\|}^{2}}}]}}$
$\leq{\vartheta_{3}}+2\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}+4G_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{z_{i,t+1}}-{x_{i,t}}}\|}^{2}}}]}}$
$\;\;+4G_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{\hat{z}}_{i,t+1}}-{z_{i,t+1}}}\|}^{2}}}]}}$
$\leq{\vartheta_{3}}+2\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}+4G_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{z_{i,t+1}}-{x_{i,t}}}\|}^{2}}}]}}$
$\;\;+4n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{t=1}^{T}{s_{t}^{2}},$ (69)
where the second inequality holds due to (46d); and the last inequality holds since (45) holds and $\{{{s_{t}}}\}$ is nonincreasing.
From ${g_{i,t}}(y)\leq{0_{{m_{i}}}}$, $\forall i\in[n]$, $\forall t\in{\mathbb{N}_{+}}$ when $y\in{\mathcal{X}_{T}}$, we have
${\Lambda_{T}}(y)\leq 0.$ (70)
Combining (69)–(70) and (46b) yields
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}\leq{\vartheta_{3}}+{\vartheta_{4}}{{\tilde{\Lambda}}_{T}}(y)+4n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}\sum\limits_{t=1}^{T}{s_{t}^{2}},\forall y\in{\mathcal{X}_{T}}.$ (71)
Using the Hölder’s inequality, we have
${\Big{(}{\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}}\Big{)}^{2}}\leq\frac{T}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}^{2}}}]}}.$ (72)
Combining (71)–(72) yields (62b).
(iii) We have
$\frac{1}{n}\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{j,t}}})}]}_{+}}}\|}]}}$ $\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}]}}}$
$=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}+{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}-{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|+\|{{g_{i,t}}({{x_{i,t}}})-{g_{i,t}}({{x_{j,t}}})}\|}]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|+{G_{2}}\|{{x_{i,t}}-{x_{j,t}}}\|}]}}},$ (73)
where the first inequality holds due to ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$; the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
From (46c), we have
$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
$\leq n{\vartheta_{1}}+{{\tilde{\vartheta}}_{2}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}$
$\leq n{\vartheta_{1}}+{{\tilde{\vartheta}}_{2}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{x_{i,t}}}\|}]}}+{{\tilde{\vartheta}}_{2}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{\hat{z}}_{i,t+1}}-{z_{i,t+1}}}\|}]}}$
$\leq n{\vartheta_{1}}+{{{\tilde{\vartheta}}_{2}}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{({{G_{1}}{\alpha_{t}}+{G_{2}}{\gamma_{0}}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]})}}+n{\tilde{p}}\sqrt{C}{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}},$ (74)
where the last inequality holds since (46e) and (45) hold, and $\{{{s_{t}}}\}$ is nonincreasing.
Combining (73)–(74) yields (62c). ∎
B. Proof of Theorem 1
(i) From (17), for any $T\in\mathbb{N}_{+}$, we have
[…]
Similar to the way to get (75), from (17) with ${\theta_{2}}\in({{\theta_{1}},1})$, for any $T\in\mathbb{N}_{+}$, we have
$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}}}}}}\leq\frac{{{{s_{0}}T^{1-{\theta_{2}}}}}}{{1-{\theta_{2}}}},$ (76)
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}-{\theta_{1}}}}}}}\leq\frac{{{{s_{0}}T^{1+{\theta_{1}}-{\theta_{2}}}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}.$ (77)
We have
$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\frac{1}{{2{\alpha_{t}}}}({{{\|{y-{x_{i,t}}}\|}^{2}}-{{\|{y-{x_{i,t+1}}}\|}^{2}}})}\Big{]}}},$
$=\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}\Big{[}{\frac{{{{\|{y-{x_{i,t}}}\|}^{2}}}}{{{\alpha_{t-1}}}}-\frac{{{{\|{y-{x_{i,t+1}}}\|}^{2}}}}{{{\alpha_{t}}}}+({\frac{1}{{{\alpha_{t}}}}-\frac{1}{{{\alpha_{t-1}}}}}){{\|{y-{x_{i,t}}}\|}^{2}}}\Big{]}}}$
$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}\Big{[}{\frac{{{{\|{y-{x_{i,1}}}\|}^{2}}}}{{{\alpha_{0}}}}-\frac{{{{\|{y-{x_{i,T+1}}}\|}^{2}}}}{{{\alpha_{T}}}}+4R{{(\mathbb{X})}^{2}}({\frac{1}{{{\alpha_{T}}}}-\frac{1}{{{\alpha_{0}}}}})}\Big{]}}$
$=\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{T}}}}\leq\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}},\forall{y}\in\mathbb{X},$ (78)
where the first inequality holds since (3) holds and $\{{{\alpha_{t}}}\}$ is nonincreasing; the last equality holds due to (3); and the last inequality holds due to (17).
Combining (62a) and (75)–(78), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}}}{{1-{\theta_{2}}}}{T^{1-{\theta_{2}}}}$
$\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}{T^{1+{\theta_{1}}-{\theta_{2}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (79)
From (17) with ${\theta_{2}}=1$, for any $T\in\mathbb{N}_{+}$, we have
[…]
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{1-{\theta_{1}}}}}}}\leq\frac{{{{s_{0}}T^{{\theta_{1}}}}}}{{{\theta_{1}}}{\alpha_{0}}}.$ (81)
Combining (62a), (75), (78) and (80)–(81), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+4n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\log(T)$
$\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{{\theta_{1}}}{\alpha_{0}}}{T^{{\theta_{1}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (82)
From (17) with ${\theta_{2}}\in({1,1+{\theta_{1}}})$, for any $T\in\mathbb{N}_{+}$, there exists a constant ${\rm Z_{1}}>0$ such that
$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}}}}}}\leq{\rm Z_{1}}{s_{0}},$ (83)
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}-{\theta_{1}}}}}}}\leq\frac{{{{s_{0}}T^{1+{\theta_{1}}-{\theta_{2}}}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}.$ (84)
Combining (62a), (75), (78) and (83)–(84), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{1}}{s_{0}}$
$\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}{T^{1+{\theta_{1}}-{\theta_{2}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (85)
From (17) with ${\theta_{2}}=1+{\theta_{1}}$, for any $T\in\mathbb{N}_{+}$, there exists a constant ${\rm Z_{2}}>0$ such that
$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}}}}}}\leq{\rm Z_{2}}{s_{0}},$ (86)
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{t}}\leq\frac{{2{s_{0}}}}{{{\alpha_{0}}}}\log(T),{\rm{if}}\;T\geq 3.$ (87)
Combining (62a), (75), (78) and (86)–(87), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{2}}{s_{0}}$
$\;\;+\frac{{4{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{{\alpha_{0}}}}\log(T)+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (88)
From (17) with ${\theta_{2}}>1+{\theta_{1}}$, for any $T\in\mathbb{N}_{+}$, there exists a constant ${\rm Z_{3}}>0$ such that
$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}}}}}}\leq{\rm Z_{3}}{s_{0}},$ (89)
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{2}}-{\theta_{1}}}}}}}\leq{\rm Z_{3}}\frac{{{s_{0}}}}{{{\alpha_{0}}}}.$ (90)
Combining (62a), (75), (78) and (89)–(90), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{3}}{s_{0}}$
$\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){\rm Z_{3}}{s_{0}}}}{{{\alpha_{0}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (91)
From (79), (82), (85), (88), and (91), we know that (20) holds.
(ii) From (17), for any $T\in\mathbb{N}_{+}$, we have
$\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}=}\sum\limits_{t=1}^{T}{\frac{{{\alpha_{t}}}}{{{\gamma_{0}}}}}\leq\frac{{{\alpha_{0}}{T^{1-{\theta_{1}}}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}},$ (92)
$\sum\limits_{t=1}^{T}\frac{{{\gamma_{0}}}}{{\gamma_{t}^{2}}}\leq{\alpha_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}}=\frac{{{\alpha_{0}^{2}}}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}}\leq\frac{{{\alpha_{0}^{2}}{T^{1-{\theta_{1}}}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}},$ (93)
$\sum\limits_{t=1}^{T}{s_{t}^{2}}\leq{s_{0}}\sum\limits_{t=1}^{T}{{s_{t}}}.$ (94)
[…]
$\sum\limits_{i=1}^{n}{\frac{{{{\|{y-{x_{i,1}}}\|}^{2}}}}{{{\gamma_{0}}}}}\leq\frac{{4nR{{(\mathbb{X})}^{2}}}}{{{\gamma_{0}}}},\forall{y}\in\mathbb{X}.$ (95)
Combining (62b) and (92)–(95), from (17) with ${\theta_{2}}\in({{\theta_{1}},1})$ and (76), we have
${({\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}})^{2}}$ $\leq{\vartheta_{3}}T+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{4}}}}{{{\gamma_{0}}}}T+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{4}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{4}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}$
$\;\;+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{4}}{s_{0}}}}{{({1-{\theta_{2}}}){\gamma_{0}}}}{T^{2-{\theta_{2}}}}+\frac{{4n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}{s_{0}^{2}}}}{{({1-{\theta_{2}}})}}{T^{2-{\theta_{2}}}}.$ (96)
Combining (62b) and (92)–(95), from (17) with ${\theta_{2}}=1$ and (80), we have
${({\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}})^{2}}$ $\leq{\vartheta_{3}}T+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{4}}}}{{{\gamma_{0}}}}T+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{4}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{4}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}$
$\;\;+\frac{{4n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{4}}{s_{0}}}}{{{\gamma_{0}}}}T\log(T)+8n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}{s_{0}^{2}}T\log(T).$ (97)
Combining (62b) and (92)–(95), from (17) with ${\theta_{2}}>1$, (83), (86), and (89), choosing ${\rm Z}=\max\{{{{\rm Z}_{1}},{{\rm Z}_{2}},{{\rm Z}_{3}}}\}$, we have
${({\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}})^{2}}$ $\leq{\vartheta_{3}}T+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{4}}}}{{{\gamma_{0}}}}T+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{4}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{4}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\gamma_{0}}}}{T^{2-{\theta_{1}}}}$
$\;\;+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{4}}{\rm Z}{s_{0}}}}{{{\gamma_{0}}}}T+4n{\tilde{p}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}{\rm Z}{s_{0}^{2}}T.$ (98)
From (96)–(98), we know that (21) holds.
(iii) We have
${\mathbf{E}_{\mathcal{C}}}[{\Lambda_{T}}({{x_{s}}})]$ $={\mathbf{E}_{\mathcal{C}}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{v_{i,t+1}^{T}{g_{i,t}}({{x_{s}}})}}{{{\gamma_{t}}}}}}\Big{]}={\mathbf{E}_{\mathcal{C}}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{[{{g_{i,t}}({{x_{i,t}}})}]_{+}^{T}{g_{i,t}}({{x_{s}}})}}\Big{]}$
$\leq-{\mathbf{E}_{\mathcal{C}}}\Big{[}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\varsigma_{s}}[{{g_{i,t}}({{x_{i,t}}})}]_{+}^{T}{\mathbf{1}_{{m_{i}}}}}}\Big{]}=-{\varsigma_{s}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}_{1}}}]}$
$\leq-{\varsigma_{s}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]},$ (99)
where the second equality holds due to (16b); and the first inequality holds due to Assumption 6.
Selecting $y={x_{s}}$ in (46b), from (16b) and (99), we have
${\varsigma_{s}}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}\leq{{\tilde{\Lambda}}_{T}}({{x_{s}}}),$ (100)
Combining (100) and (92)–(95), from (17) with ${\theta_{2}}\in({{\theta_{1}},1})$ and (76), we have
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}$ $\leq\frac{{2nR{{(\mathbb{X})}^{2}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{({1-{\theta_{2}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{2}}}}.$ (101)
Combining (62c), (75), (101), from (17) with ${\theta_{2}}\in({{\theta_{1}},1})$ and (76), we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $\leq n{G_{2}}{\vartheta_{1}}+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{6}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{6}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{6}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{{\vartheta_{5}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{6}}{s_{0}}}}{{({1-{\theta_{2}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{2}}}}+\frac{{n{\tilde{p}}\sqrt{C}{G_{2}}{{\tilde{\vartheta}}_{2}}{s_{0}}}}{{({1-{\theta_{2}}})}}{T^{1-{\theta_{2}}}}.$ (102)
Combining (100) and (92)–(95), from (17) with ${\theta_{2}}=1$ and (80), we have
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}$ $\leq\frac{{2nR{{(\mathbb{X})}^{2}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{4n{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{{\varsigma_{s}}{\gamma_{0}}}}\log(T).$ (103)
Combining (62c), (75), (103), from (17) with ${\theta_{2}}=1$ and (80), we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $\leq n{G_{2}}{\vartheta_{1}}+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{6}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{6}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{6}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{{\vartheta_{5}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{4n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{6}}{s_{0}}}}{{{\varsigma_{s}}{\gamma_{0}}}}\log(T)+2n{\tilde{p}}\sqrt{C}{G_{2}}{{\tilde{\vartheta}}_{2}}{s_{0}}\log(T).$ (104)
Combining (100) and (92)–(95), from (17) with ${\theta_{2}}>1$, (83), (86), and (89), we have
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}$ $\leq\frac{{2nR{{(\mathbb{X})}^{2}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\rm Z}{s_{0}}}}{{{\varsigma_{s}}{\gamma_{0}}}}.$ (105)
Combining (62c), (75), (105), from (17) with ${\theta_{2}}>1$, (83), (86), and (89), we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $\leq n{G_{2}}{\vartheta_{1}}+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{6}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{2n{G_{1}}R(\mathbb{X}){\vartheta_{6}}{\alpha_{0}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}+\frac{{2nG_{1}^{2}{\vartheta_{6}}{\alpha_{0}^{2}}}}{{({1-{\theta_{1}}}){\varsigma_{s}}{\gamma_{0}}}}{T^{1-{\theta_{1}}}}$
$\;\;+\frac{{{\vartheta_{5}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X}){\vartheta_{6}}{\rm Z}{s_{0}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+n{\tilde{p}}\sqrt{C}{G_{2}}{{\tilde{\vartheta}}_{2}}{\rm Z}{s_{0}}.$ (106)
From (102), (104), and (106), we know that (22) holds.
C. Proof of Theorem 2
(i) From (23), for any $T\in\mathbb{N}_{+}$, we have
${\alpha_{T}}={\alpha_{0}}\sqrt{\frac{{{\Psi_{T}}}}{T}}\leq\frac{{{\alpha_{0}}\mu}}{{1-\mu}},$ (107)
$\sum\limits_{t=1}^{T}{{\alpha_{t}}}={\alpha_{0}}\sum\limits_{t=1}^{T}{\sqrt{\frac{{{\Psi_{t}}}}{t}}\leq{\alpha_{0}}\sqrt{{\Psi_{T}}}}\sum\limits_{t=1}^{T}{\frac{1}{{\sqrt{t}}}}\leq 2{\alpha_{0}}\sqrt{{\Psi_{T}}T}\leq\frac{{2{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T},$ (108)
$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{{\mu^{t}}}=\frac{{{s_{0}}\mu({1-{\mu^{T}}})}}{{1-\mu}}\leq\frac{{{s_{0}}\mu}}{{1-\mu}},$ (109)
$\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{{{\mu^{t}}\sqrt{t}}}{{\sqrt{{\Psi_{t}}}}}}\leq\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\sqrt{{\mu^{t}}}}\leq\frac{{{s_{0}}\sqrt{\mu}}}{{{\alpha_{0}}({1-\sqrt{\mu}})}}.$ (110)
Combining (62a), (78), (107)–(110), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{2{\vartheta_{2}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\mu}}{{1-\mu}}$
$\;\;+\frac{{2\tilde{p}\sqrt{C}R(\mathbb{X}){s_{0}}\sqrt{\mu}}}{{{\alpha_{0}}({1-\sqrt{\mu}})}}+\frac{{2R{{(\mathbb{X})}^{2}}({1-\mu})}}{{{\alpha_{0}}\mu}}.$ (111)
From (111), we know that (24) holds.
(ii) From (17), for any $T\in\mathbb{N}_{+}$, we have
$\sum\limits_{t=1}^{T}{\frac{1}{{{\gamma_{t}}}}}=\frac{1}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{{\alpha_{t}}\leq}\frac{{2{\alpha_{0}}\mu}}{{{\gamma_{0}}({1-\mu})}}\sqrt{T},$ (112)
$\sum\limits_{t=1}^{T}{\frac{{{\gamma_{0}}}}{{\gamma_{t}^{2}}}}=\frac{{\alpha_{0}^{2}}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\frac{{{\Psi_{t}}}}{t}}\leq\frac{{\alpha_{0}^{2}{\Psi_{T}}}}{{{\gamma_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{t}}\leq\frac{{2\alpha_{0}^{2}\mu}}{{{\gamma_{0}}({1-\mu})}}\log(T),{\rm{if}}\;T\geq 3,$ (113)
$\sum\limits_{t=1}^{T}{s_{t}^{2}}=s_{0}^{2}\sum\limits_{t=1}^{T}{{\mu^{2t}}}\leq\frac{{s_{0}^{2}{\mu^{2}}}}{{1-{\mu^{2}}}}.$ (114)
Combining (62b), (95), (109), and (112)–(114), we have
${({\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{t}}({{x_{i,t}}})}]}_{+}}}\|}]}}})^{2}}$ $\leq{\vartheta_{3}}T+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{4}}}}{{{\gamma_{0}}}}T+\frac{{4n{G_{1}}R(\mathbb{X}){\vartheta_{4}}{\alpha_{0}}\mu}}{{{\gamma_{0}}({1-\mu})}}{T^{3/2}}+\frac{{4nG_{1}^{2}{\vartheta_{4}}\alpha_{0}^{2}\mu}}{{{\gamma_{0}}({1-\mu})}}T\log(T)$
$\;\;+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X}){\vartheta_{4}}{s_{0}}\mu}}{{{\gamma_{0}}({1-\mu})}}T+\frac{{4n{{\tilde{p}}^{2}}CG_{2}^{2}{{\tilde{\vartheta}}_{4}}s_{0}^{2}{\mu^{2}}}}{{1-{\mu^{2}}}}T.$ (115)
From (115), we know that (25) holds.
(iii) Combining (100), (95), (109), and (112)–(113), we have
$\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|}]}}$ $\leq\frac{{2nR{{(\mathbb{X})}^{2}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{4n{G_{1}}R(\mathbb{X}){\alpha_{0}}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}\sqrt{T}+\frac{{4nG_{1}^{2}\alpha_{0}^{2}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}\log(T)$
$\;\;+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X}){s_{0}}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}.$ (116)
Combining (62c), (108), and (109), we have
$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $\leq n{G_{2}}{\vartheta_{1}}+\frac{{2{\vartheta_{5}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{2nR{{(\mathbb{X})}^{2}}{\vartheta_{6}}}}{{{\varsigma_{s}}{\gamma_{0}}}}+\frac{{4n{G_{1}}R(\mathbb{X}){\vartheta_{6}}{\alpha_{0}}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}\sqrt{T}$
$\;\;+\frac{{4nG_{1}^{2}{\vartheta_{6}}\alpha_{0}^{2}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}\log(T)+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X}){\vartheta_{6}}{s_{0}}\mu}}{{{\varsigma_{s}}{\gamma_{0}}({1-\mu})}}+\frac{{n\tilde{p}\sqrt{C}{G_{2}}{{\tilde{\vartheta}}_{2}}{s_{0}}\mu}}{{1-\mu}}.$ (117)
From (117), we know that (26) holds.

### v2

@@
 We show that the proposed algorithm establishes an $\mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}})$ network regret bound and an $\mathcal{O}({T^{1-{\theta_{1}}/2}})$ network cumulative constraint violation bound, where $T$ is the number of iterations and ${\theta_{1}}\in({0,1})$ is a user-defined trade-off parameter.
-When Slater’s condition holds (i.e, there is a point that strictly satisfies the inequality constraints at all iterations), the network cumulative constraint violation bound is reduced to $\mathcal{O}({T^{1-{\theta_{1}}}})$.
+When Slater’s condition holds, the network cumulative constraint violation bound is reduced to $\mathcal{O}({T^{1-{\theta_{1}}}})$.
 These bounds are comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication for distributed online convex optimization with (time-varying) inequality constraints.
@@
 Distributed online convex optimization offers a promising framework for modeling a variety of problems in dynamic, uncertain, and adversarial environments, with wide-ranging applications such as real-time routing in data networks and online advertisement placement in web search [1].
-This framework can be understood as a structured repeated game with $T$ iterations between a network of agents and an adversary.
-Specifically, at each iteration $t$, each agent $i$ selects a decision ${x_{i,t}}\in\mathbb{X}$, where $\mathbb{X}\subseteq{\mathbb{R}^{p}}$ is a known convex set and $p$ is a positive integer.
-Upon selection, the local loss function ${f_{i,t}}$ is privately revealed to agent $i$ by the adversary.
-The goal of agents is to collaboratively minimize the network-wide accumulated loss, and the corresponding performance metric is network regret
-$\frac{1}{n}\sum\limits_{i=1}^{n}{\Big{(}{\sum\limits_{t=1}^{T}{{f_{t}}({{x_{i,t}}})}-\mathop{\min}\limits_{x\in\mathbb{X}}\sum\limits_{t=1}^{T}{{f_{t}}({{x}})}}\Big{)}},$
-where ${f_{t}}(x):=\frac{1}{n}\sum\nolimits_{j=1}^{n}{{f_{j,t}}(x)}$ is the global loss function of the network at iteration $t$.
 Various projection-based distributed online algorithms with sublinear regret have been proposed to solve the distributed online convex optimization problem, see, e.g., [2, 3, 4, 5, 6, 7, 8, 9, 10], and recent survey paper [11].
@@
 The aforementioned distributed online algorithms rely on agents exchanging their local data with perfect communication.
-Consequently, these algorithms encounter significant limitations arising from communication bottlenecks, particularly in scenarios involving limited bandwidth and power resources.
-The limitations becomes more pronounced as the scale of the multi-agent network and the dimensionality of the exchanged data increase.
+Consequently, these algorithms encounter significant limitations arising from communication bottlenecks.
 To overcome the limitations, distributed online algorithms with compressed communication are studied for the fix communication topology in the literature, see [12, 13, 14], and recent survey paper [15].
@@
 $\mathbb{N}_{+}$}, $\mathbb{R}$, $\mathbb{R}^{p}$} and $\mathbb{R}_{+}^{p}$ denote the sets of all positive integers, real numbers, $p$-dimensional and nonnegative vectors, respectively.
+Let $\mathbb{X}$ denote a subset of $\mathbb{R}^{p}$}.
 Given $m$ and $n\in{\mathbb{N}_{+}}$, $[m]$ denotes the set $\{{1,\cdot\cdot\cdot,m}\}$, and $[m,n]$ denotes the set $\{{m,\cdot\cdot\cdot,n}\}$ for $m<n$.
 Given vectors $x$ and $y$, $x^{T}$} denotes the transpose of the vector $x$, and $\langle{x,y}\rangle$ denotes the standard inner.
-${\mathbf{0}_{p}}$ and ${\mathbf{1}_{p}}$ denote the $p$-dimensional column vector whose components are all 00 and $1$, respectively.
+$\mathbf{0}_{p}$} and $\mathbf{1}_{p}$} denote the $p$-dimensional column vector whose components are all 0 and $1$, respectively.
 $\mathrm{col}({q_{1}},\cdot\cdot\cdot,{q_{n}})$ denotes the concatenated column vector of ${q_{i}}\in{\mathbb{R}^{{m_{i}}}}$ for $i\in[n]$.
 $\mathbb{B}^{p}$} and $\mathbb{S}^{p}$} denote the unit ball and sphere centered around the origin in $\mathbb{R}^{p}$} under Euclidean norm, respectively.
-For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $[x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$.
+For a set $\mathbb{K}\in{\mathbb{R}^{p}}$ and a vector $x\in{\mathbb{R}^{p}}$, ${\mathcal{P}_{\mathbb{K}}}(x)$ denotes the projection of the vector $x$ onto the set $\mathbb{K}$, i.e., ${\mathcal{P}_{\mathbb{K}}}(x)=\arg{\min_{y\in{\mathbb{K}}}}{\|{x-y}\|^{2}}$, and $x]_{+}$ denotes ${\mathcal{P}_{\mathbb{R}_{+}^{p}}}(x)$.
 For a function $f$ and a vector $x$, $\nabla f(x)$ denotes the subgradient of $f$ at $x$.
@@
 Assumption 3.
-For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant ${L}$ such that
+For all $i\in[n]$, $t\in{\mathbb{N}_{+}}$, there exists a constant L{L} such that
 $\|{\nabla{f_{i,t}}(x)-\nabla{f_{i,t}}(y)}\|\leq{L}\|{x-y}\|,x,y\in\mathbb{X}.$ (6)
-Remark 1.
-Note that we do not require the assumptions that the local loss function ${{f_{i,t}}}$ and local constraint function ${{g_{i,t}}}$ are uniformly bounded, whereas [20] imposes them, and [24] imposes the first one.
-In addition, Assumption 2 is commonly used for distributed online convex optimization with long-term constraints [16, 18, 20, 24].
-Since the local constraint function ${{g_{i,t}}}$ is convex, Assumption 2 implies its Lipschitz continuity on $\mathbb{X}$.
-In contrast, the nonconvex nature of the local constraint function ${{f_{i,t}}}$ precludes the existence of such a Lipschitz continuity property.
-This fundamental distinction causes analytical challenges compared to distributed online convex optimization, particularly in establishing network regret bounds.
-Furthermore, Assumption 3 plays a crucial role in addressing the analytical challenges, which is also used in distributed online nonconvex optimization [25].
-Different from the studies on distributed online convex optimization with compressed communication [27, 12, 13, 14] that consider a fix and undirected graph, we consider a time-varying directed graph.
-The following assumption on the graph is made, which is also used in [16, 20, 24].
 Assumption 4.
@@
 (iii) There exists an integer $B>0$ such that the time-varying directed graph $({\mathcal{V},\cup_{l=0}^{B-1}{\mathcal{E}_{t+l}}})$ is strongly connected.
-In this paper, we consider the scenario where the communication between agents is compressed by using a class of compressors with globally bounded absolute compression error, as given in the following.
 Assumption 5.
-The compressor $\mathcal{C}:{\mathbb{Y}}\to{\mathbb{Y}}$ with $\mathbb{Y}\in{\mathbb{R}^{p}}$ satisfies
-${\mathbf{E}_{\mathcal{C}}}[{\|{\mathcal{C}(x)-x}\|_{d}^{2}}]\leq C,\forall x\in{\mathbb{Y}},$ (7)
+The compressor $\mathcal{C}:{\mathbb{R}^{p}}\to{\mathbb{R}^{p}}$ satisfies
+${\mathbf{E}_{\mathcal{C}}}[{\|{\mathcal{C}(x)-x}\|_{d}^{2}}]\leq C,\forall x\in{\mathbb{R}^{p}},$ (7)
 for some real norm parameter $d\geq 1$ and constant $C\geq 0$.
 Here $\mathbf{E}_{\mathcal{C}}$} denotes the expectation over the internal randomness of the stochastic compression operator $\mathcal{C}$.
-Assumption 5 covers the majority of popular compressors in machine learning and signal processing applications such as the deterministic quantization used in [28, 29] and the unbiased stochastic quantization used in [30, 31].
-The same class of compressors satisfying Assumption 5 is also used in [32, 14, 33].
-Next, Slater’s condition is present in the following.
 Assumption 6.
@@
 ${g_{t}}({{x_{s}}})\leq-{\varsigma_{s}}{\mathbf{1}_{m}},t\in{\mathbb{N}_{+}}.$ (8)
-Slater’s condition is a sufficient condition for strong duality to hold in convex optimization problems [21].
-However, for nonconvex problems, Slater’s condition alone does not generally guarantee strong duality because the standard Lagrange dual often has a nonzero duality gap in nonconvex problems [34].
-For distributed online convex optimization, [24] establishes reduced network cumulative constraint violation bounds under Slater’s condition.
-However, to the best of our knowledge, there are no studies to show that similar reductions for cumulative constraint violation bounds can be achieved in distributed online nonconvex optimization, even for the general constraint violation bounds.
 III Distributed Online Primal–Dual Algorithm with
 Compressed Communication
-In this section, we propose a distributed online primal–dual algorithm with compressed communication by using the class of compressors satisfying Assumption 5, and analyze the performance of the algorithm without and with Slater’s condition, respectively.
 III-A Algorithm Description
@@
 where $\gamma_{t}$}} is the regularization parameter; and $\alpha_{t}$}} is the stepsize.
-To implement the algorithm (9), at each iteration each agent $i$ need to receive the exact vector-valued variable $z_{j,t}$ from its neighbors.
-That results in a substantial amount of data exchange over all iterations, especially when the scale of the multi-agent network and the dimension $p$ are large.
-To improve communication efficiency, we use the compressor satisfying Assumption 5 to compress $z_{j,t}$.
-Note that directly using the compressed variable $\mathcal{C}({{z_{j,t}}})$ to replace $z_{j,t}$ in (9a) does not work due to the compression error.
-To reduce the compression error, an auxiliary variable $\hat{z}_{j,t}\in{\mathbb{R}^{p}}$ and a scaling parameter ${s_{t}}$ can be introduced.
-Instead of directly compressing $z_{j,t}$, one can compress the scaled difference ${({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}$, then multiply the compressed result by ${s_{t}}$ and add it back to ${{\hat{z}}_{j,t-1}}$ to replace $z_{j,t}$ in (9a).
-Thus, (9a) is modified as follows:
-${x_{i,t}}=\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\hat{z}_{j,t}}},$ (10)
-where
-${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)}.$ (11)
-In this way, the compression error is reduced.
-However, as stated in [32], at each iteration each agent $i$ needs to receive the exact vector-valued variable $\hat{z}_{j,t-1}$ from its neighbors due to the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$ in (10).
-Therefore, the improved algorithm does not enjoy the benefits of compression.
-To deal with the dilemma, [32] introduces another auxiliary variable ${y_{j,t}}\in{\mathbb{R}^{p}}$ to remove the term $\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t-1}}}$.
-Then, (9a) becomes the following form:
-${x_{i,t}}={{\hat{z}}_{j,t}}-{y_{j,t}},$ (12)
-where
-${{\hat{z}}_{j,t}}$ $={{\hat{z}}_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)},$ (13a)
-${y_{j,t}}$ $={y_{j,t-1}}+{s_{t}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}-{s_{t}}\sum\limits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{\cal C}\big{(}({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}}\big{)}}.$ (13b)
-Note that agent $i$ only requires the compressed data as presented in (13).
-The similar strategy is also used in [13, 14].
-When the graph $\mathcal{G}_{t}$ is fixed, i.e., ${[{W_{t-1}}]_{ij}}={[{W_{t}}]_{ij}}$ for all $t\in[T]$, it is straightforward to check that ${y_{j,t}}={{\hat{z}}_{j,t}}-\sum\nolimits_{j=1}^{n}{{{[{W_{t}}]}_{ij}}{{\hat{z}}_{j,t}}}$ for $j\in[n]$ by mathematical induction.
-However, the critical result does not hold when the graph $\mathcal{G}_{t}$ is time-varying.
-To handle this challenge, instead of introducing another auxiliary variable, we define $\hat{z}_{j,t}$ in (10) as a vector-valued variable stored in agent $i$ rather than its neighbors.
-The variable $\hat{z}_{j,t}$ is indexed by $j$ to maintain one-to-one correspondence with all agents.
-Consequently, its value no longer needs to be exchanged, and only the compressed data $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$ in (11) is required by agent $i$.
-In this way, the distributed online primal–dual algorithm with compressed communication is proposed, which is presented in pseudo-code as Algorithm 1.
+To improve communication efficiency, the distributed online primal–dual algorithm with compressed communication is proposed by using the class of compressors satisfying Assumption 5, which is presented in pseudo-code as Algorithm 1.
 Algorithm 1 Distributed Online Primal–Dual Algorithm with Compressed Communication
 0: non-increasing stepsize sequence $\{{\alpha_{t}}\}\subseteq({0,+\infty})$, non-increasing scaling parameter sequence $\{{{s_{t}}}\}\subseteq({0,+\infty})$, and non-decreasing regularization parameter sequence $\{{\gamma_{t}}\}\subseteq({0,+\infty})$.
-0:  ${\hat{z}_{j,0}}={\mathbf{0}_{p}}$ for $j\in[n]$, and ${z_{i,1}}\in{\mathbb{X}}$, $\forall i\in[n]$.
+[…]
 for $t=1,\cdot\cdot\cdot$ do
@@
    Update
-${\hat{z}_{j,t}}$ $={\hat{z}_{j,t-1}}+{s_{t}}\mathcal{C}\big{(}{({{z_{j,t}}-{\hat{z}_{j,t-1}}})/{s_{t}}}\big{)},j\in\big{\{}\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})\cup\{i\}\big{\}}.$ (14)
+${{\hat{z}}_{j,t}}={\mathcal{P}_{\mathbb{X}}}\big{(}{{{\hat{z}}_{j,t-1}}+{s_{t}}{\cal C}(({z_{j,t}}-{{\hat{z}}_{j,t-1}})/{s_{t}})}\big{)},j\in\big{\{}\mathcal{N}_{i}^{\text{in}}({{\mathcal{G}_{t}}})\cup\{i\}\big{\}}.$ (10)
    Select
@@
 $\{x_{i,t}\}$.
-It is worth noting that agent $j$ does not need to receive ${{\hat{z}}_{j,t-1}}$ in (14) from any other agents to compute $\mathcal{C}({({{z_{j,t}}-{{\hat{z}}_{j,t-1}}})/{s_{t}}})$, as it naturally possesses this information.
-In Algorithm 1, we initialize ${\hat{z}_{j,0}}={\mathbf{0}_{p}}$ for $j\in[n]$, and thus ${\hat{z}_{j,1}}={s_{1}}{\cal C}({z_{j,1}}/{s_{1}})$, which is inherently known by agent $j$ since $z_{j,1}$ is precisely stored in agent $j$.
-More generally, at each iteration $t$, agent $j$ has access to both $z_{j,t}$ and $\hat{z}_{j,t-1}$, enabling it to independently determine $\hat{z}_{j,t}$ without requiring information from any other agents.
 III-B Performance Analysis
@@
 where ${\theta_{1}}\in({0,1})$, ${\alpha_{0}}>0$, ${\gamma_{0}}\in({0,1/({4G_{2}^{2}})}]$, ${s_{0}}>0$, and ${\theta_{2}}>{\theta_{1}}$ are constants. Then, for any $T\in{\mathbb{N}_{+}}$,
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]=\left\{\begin{array}[]{l}\mathcal{O}({{T^{\max\{{1-{\theta_{1}},1+{\theta_{1}}-{\theta_{2}}}\}}}}),{\rm{if}}\;{\theta_{1}}<{\theta_{2}}<1,\\ \mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}}),{\rm{if}}\;{\theta_{2}}\geq 1,\end{array}\right.$ (20)
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]=\left\{\begin{array}[]{l}\mathcal{O}({{T^{\max\{{1-{\theta_{1}},1+{\theta_{1}}-{\theta_{2}}}\}}}}),{\rm{if}}\;{\theta_{1}}<{\theta_{2}}<1,\\ \mathcal{O}({{T^{\max\{{1-{\theta_{1}},{\theta_{1}}}\}}}}),{\rm{if}}\;{\theta_{2}}\geq 1,\end{array}\right$.
+(16)
 $\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]=\mathcal{O}({T^{1-{\theta_{1}}/2}}).$ (17)
@@
 $\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}({T^{1-{\theta_{1}}}}).$ (18)
-Due to the space limitations, we omit the proof, which can be found in the arXiv version
 Remark 1.
@@
 These bounds characterize the impact of compressed communication on the network regret and cumulative constraint violation bounds, which is captured by $\theta_{2}$}.
-When ${\theta_{2}}\geq 1$, compressed communication does not affect either of the bounds, and they are the same as the state-of-the-art results established by the distributed online algorithms without compressed communication in [20, 24].
-When ${\theta_{1}}<{\theta_{2}}<1$, compressed communication may enable the network regret bound to become larger due to $01-{\theta_{2}}>$ 0, but does not affect the network cumulative constraint violation bound due to ${\theta_{2}}>{\theta_{1}}$.
+When ${\theta_{2}}\geq 1$, they are the same as the state-of-the-art results established by the distributed online algorithms without compressed communication in [20, 24].
 In addition, when Slater’s condition holds, the network cumulative constraint violation bound is further reduced as in (18).
-The bound is unaffected by compressed communication and remains the same as the results established by the distributed online algorithm with perfect communication in [24].
+The bound remains the same as the results established by the distributed online algorithm with perfect communication in [24].
 We then choose the scaling parameter sequence $\{s_{t}\}$ produced by $\{\mu^{t}\}$, which is also adopted by the distributed algorithms in [27, 14].
@@
 $\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{CCV}}(T)}]$ $=\mathcal{O}(\sqrt{T}).$ (22)
-Due to the space limitations, we omit the proof, which can be found in the arXiv version
 Remark 2.
@@
 $\mathop{\min}\limits_{x}\;\;\;\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{\frac{1}{4}}}{\big{|}{{{\|{{S_{i}}-x}\|}^{2}}-{D_{i,t}}}\big{|}^{2}},$ (23a)
-$\;{\rm{s}}{\rm{.t}}{\rm{.}}\;\;\;\;x\in\mathbb{X},{\rm{}}{B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in[T].$ (27b)
-In this problem, the sensors aim to cooperatively track a moving target.
-The position of sensor $i$ is denoted by $S_{i}\in{\mathbb{R}^{p}}$.
-Sensor $i$ measures the distance between the positions of the target and itself by ${D_{i,t}}={\|{{S_{i}}-{X_{0,t}}}\|^{2}}+{\tau_{i,t}}$ where ${X_{0,t}}\in{\mathbb{R}^{p}}$ and ${\tau_{i,t}}\in\mathbb{R}$ denote the position of the target and the measurement noise at iteration $t$, respectively.
-Moreover, each agent only has access to its distance measurement.
-In addition, the position of the target is required to remain within a designated safe region characterized by the set $\mathbb{X}$, and avoid prolonged deviations from the mission region characterized by time-varying linear inequality constraint ${B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}}$ with ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ and ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$.
-However, occasional deviations are permitted to accommodate obstacle avoidance and exploration requirements.
-Inspired by [35, 36, 37], the problem as in (27) is formulated to achieve the least squares estimator for the position of the target.
+$\;{\rm{s}}{\rm{.t}}{\rm{.}}\;\;\;\;x\in\mathbb{X},{\rm{}}{B_{i,t}}x-{b_{i,t}}\leq{\mathbf{0}_{{m_{i}}}},\forall i\in[n],\forall t\in[T],$ (23b)
+where $S_{i}\in{\mathbb{R}^{p}}$ denotes the position of sensor $i$, ${D_{i,t}}={\|{{S_{i}}-{X_{0,t}}}\|^{2}}+{\tau_{i,t}}$ denotes the distance between the positions of the target and sensor $i$, ${X_{0,t}}\in{\mathbb{R}^{p}}$ and $\tau_{i,t}$$\in\mathbb{R}$ denote the position of the target and the measurement noise at iteration $t$, respectively, and ${B_{i,t}}\in{\mathbb{R}^{{m_{i}}\times p}}$ and ${b_{i,t}}\in{\mathbb{R}^{{m_{i}}}}$ denote the coefficient matrix and coefficient vector of the local linear constraints, respectively.
 The communication topology is modeled by a time-varying undirected graph.
 Specifically, at each iteration $t$, the graph is first randomly generated where the probability of any two sensors being connected is $\rho$.
-Then, to make sure that Assumption 4 is satisfied, we add edges $({i,i+1})$ for $i\in[24]$ when $t\in\{{4c+1}\}$, edges $({i,i+1})$ for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges $({i,i+1})$ for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges $({i,i+1})$ for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$.
+Then, to make sure that Assumption 4 is satisfied, we add edges (i,i+1)({i,i+1}) for $i\in[24]$ when $t\in\{{4c+1}\}$, edges (i,i+1)({i,i+1}) for $i\in[25,49]$ when $t\in\{{4c+2}\}$, edges (i,i+1)({i,i+1}) for $i\in[50,74]$ when $t\in\{{4c+3}\}$, edges (i,i+1)({i,i+1}) for $i\in[75,99]$ when $t\in\{{4c+4}\}$ for $c=\{0,1,\cdot\cdot\cdot\}$.
 Moreover, let ${[{{W_{t}}}]_{ij}}=\frac{1}{n}$ if $({j,i})\in{\mathcal{E}_{t}}$ and ${[{{W_{t}}}]_{ii}}=1-\sum\nolimits_{j=1}^{n}{{{[{{W_{t}}}]}_{ij}}}$.
@@
 To verify the theoretical results, we compare Algorithm 1 with the algorithm in [24].
-We set $\rho=0.1$, $\mathbb{X}={[{-5,5}]^{p}}$, $p=2$, ${m_{i}}=2$, and randomly choose each component of ${{S_{i}}}$ from the uniform distribution in the interval $[{-10,10}]$.
+We set $\rho=0.1$, $\mathbb{X}={[{-5,5}]^{p}}$, $p=2$, ${m_{i}}=2$, and randomly choose each component of $S_{i}$}} from the uniform distribution in the interval [10,10][{-10,10}].
 We assume that the position of the target evolves by
@@
 where $Q_{t}$}} is randomly generated from Bernoulli distribution with a success probability of $0.5$, and ${X_{0,0}}={[{0.8,0.95}]^{T}}$.
-Moreover, ${\tau_{i,t}}$ is randomly generated from the uniform distribution from in the interval $[{0,0.001}]$.
-Furthermore, each component of ${B_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{0,2}]$, and each component of ${b_{i,t}}$ is randomly generated from the uniform distribution in the interval $[{b,b+1}]$ with $b>0$.
+Moreover, $\tau_{i,t}$} is randomly generated from the uniform distribution from in the interval [0,0.001][{0,0.001}].
+Furthermore, each component of $B_{i,t}$} is randomly generated from the uniform distribution in the interval [0,2][{0,2}], and each component of $b_{i,t}$} is randomly generated from the uniform distribution in the interval [b,b+1][{b,b+1}] with $b>0$.
 Note that $b>0$ guarantees Slater’s condition holds.
 Here we choose $b=0.01$.
-In addition, we select the following compressor for Algorithm 1.
-Standard uniform quantizer:
+In addition, we select the following compressor for Algorithm 1:
 $\mathcal{C}(x)=\Delta\Big{\lfloor}{\frac{x}{\Delta}+\frac{{{\mathbf{1}_{p}}}}{2}}\Big{\rfloor},$
@@
 Here we set $\Delta=1$ and $q=8$.
-The inputs of all these algorithms are listed in TABLE I.
-TABLE I: Input of algorithms.
-Algorithms Inputs
-Algorithm 1 in this paper ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},{s_{t}}=1/t$
-The algorithm in [24] ${\alpha_{t}}=0.5/t,{\gamma_{t}}=0.15/{\alpha_{t}},\psi(x)={\|x\|^{2}}/2$
 Figure 1: Evolutions of network regret.
@@
 Moreover, the network cumulative constraint violation bounds were further reduced when Slater’s condition held.
-These bounds were comparable to the state-of-the-art results established by existing distributed online algorithms with perfect communication, even in the context of distributed online convex optimization with (time-varying) inequality constraints.
 In the future, we plan to investigate distributed bandit nonconvex optimization with time-varying constraints since gradient information is unavailable in many real-world applications.
@@
 $\leq{{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{{\|{{z_{i,t+1}}-{\hat{z}_{i,t+1}}}\|_{d}^{2}}}]$
-$={{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t}}-{s_{t+1}}\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
+$\leq{{\tilde{p}}^{2}}{\mathbf{E}_{\mathcal{C}}}[{\|{{z_{i,t+1}}-{\hat{z}_{i,t}}-{s_{t+1}}\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
 $={{\tilde{p}}^{2}}s_{t+1}^{2}{\mathbf{E}_{\mathcal{C}}}[{\|{({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}-\mathcal{C}({({{z_{i,t+1}}-{\hat{z}_{i,t}}})/{s_{t+1}}})}\|_{d}^{2}}]$
@@
 ${\vartheta_{1}}=\frac{{2\tau}}{{\lambda({1-\lambda})}}\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|},{{\tilde{\vartheta}}_{2}}=\frac{{4-4\lambda+2n\tau}}{{1-\lambda}},$
-${{\tilde{\vartheta}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\vartheta}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{\left({1-\lambda}\right)}^{2}}}}+32.$
+${{\tilde{\vartheta}}_{3}}=\frac{{16n{\tau^{2}}}}{{{\lambda^{2}}({1-{\lambda^{2}}})}}{\Big{(}{\sum\limits_{i=1}^{n}{\|{{\hat{z}_{i,1}}}\|}}\Big{)}^{2}},{{\tilde{\vartheta}}_{4}}=\frac{{16{n^{2}}{\tau^{2}}}}{{{{({1-\lambda})}^{2}}}}+32.$
 Proof.
@@
 Under the same conditions as stated in Lemma 6, and supposing that Assumption 3 holds, for any $T\in{\mathbb{N}_{+}}$ and any $y\in{\mathcal{X}_{T}}$, it holds that
-$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}\leq 2nR(\mathbb{X})L{\vartheta_{1}}+{\vartheta_{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}$
+$\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}\leq\hat{\vartheta}+{\vartheta_{2}}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}\sum\limits_{t=1}^{T}{{s_{t}}}$
 $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+2{\tilde{p}}\sqrt{C}R(\mathbb{X})\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{{\Delta_{i,t}}(y)}]}},$ (58a)
@@
 where
-${\vartheta_{2}}=2G_{1}^{2}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2},{\vartheta_{3}}=2G_{2}^{2}{{\tilde{\vartheta}}_{3}},{\vartheta_{4}}=\frac{{4\max\{{1,2G_{2}^{2}{{\tilde{\vartheta}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},{\vartheta_{5}}={n{G_{1}}{G_{2}}{{\tilde{\vartheta}}_{2}}},$
-${\vartheta_{6}}={1+G_{2}^{2}{\gamma_{0}}{{\tilde{\vartheta}}_{2}}}.$
+$\hat{\vartheta}=2R(\mathbb{X})L{\vartheta_{1}}+{G_{1}}{\vartheta_{1}},{{\tilde{\vartheta}}_{1}}=2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}+{G_{1}}{{\tilde{\vartheta}}_{2}},{\vartheta_{2}}=2G_{1}^{2}+\tilde{\vartheta}_{1}^{2},{\vartheta_{3}}=2G_{2}^{2}{{\tilde{\vartheta}}_{3}},$
+${\vartheta_{4}}=\frac{{4\max\{{1,2G_{2}^{2}{{\tilde{\vartheta}}_{4}}}\}}}{{\min\{{1,\frac{1}{{2{\gamma_{0}}}}}\}}},{\vartheta_{5}}={n{G_{1}}{G_{2}}{{\tilde{\vartheta}}_{2}}},{\vartheta_{6}}={1+G_{2}^{2}{\gamma_{0}}{{\tilde{\vartheta}}_{2}}}.$
 Proof.
@@
 $=\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{i,t}}({{x_{i,t}}})}]}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\nabla{f_{j,t}}({{x_{i,t}}})-\nabla{f_{j,t}}({{x_{j,t}}})}]}}.$ (60)
-From (64) and (6), we have
+From (60), (6) and (3), we have
 $\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}$
-$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{L\|{{x_{i,t}}-{x_{j,t}}}\|\|{{x_{i,t}}-y}\|}]}}}$
-$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}},\forall y\in\mathbb{X},$ (65)
-where last inequality holds due to (3).
-For the second term on the right-hand side of (65), from (46c), we have
-$\;\;\;\;\;\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
-$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{\varepsilon_{i,t}^{z}}\|}]}}$
-$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{z_{i,t+1}}-{{\hat{z}}_{i,t+1}}}\|}]}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{{x_{i,t}}-{z_{i,t+1}}}\|}]}}$
-$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{4nR{{(\mathbb{X})}^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}{\alpha_{t}}+\frac{{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}}{{4n{\alpha_{t}}}}}\Big{]}}}$
-$\leq 2nR(\mathbb{X})L{\vartheta_{1}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\sum\limits_{t=1}^{T}{{s_{t}}}+4n^{2}R{(\mathbb{X})^{2}}{L^{2}}\tilde{\vartheta}_{2}^{2}\sum\limits_{t=1}^{T}{{\alpha_{t}}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{\frac{{{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{z_{i,t+1}}}\|}^{2}}}]}}{{4{\alpha_{t}}}}}},$ (66)
-where the second inequality holds due to $\varepsilon_{i,t}^{z}={{\hat{z}}_{i,t+1}}-{x_{i,t}}$; and the third inequality holds since (45) holds and $\{{{s_{t}}}\}$ is nonincreasing.
+$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{t}}({{x_{i,t}}}),{{\bar{z}}_{t}}-y}\rangle}]}}+\frac{1}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{\nabla{f_{j,t}}({{x_{i,t}}})}\|\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}}$
+$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{{\bar{z}}_{t}}-y}\rangle}]}}+\frac{{2R(\mathbb{X})L}}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
+$\;\;+\frac{{{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}$
+$\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\langle{\nabla{f_{i,t}}({{x_{i,t}}}),{x_{i,t}}-y}\rangle}]}}+\frac{{2R(\mathbb{X})L}}{{{n^{2}}}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}$
+$\;\;+\frac{{2{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{x_{i,t}}-{{\bar{z}}_{t}}}\|}]}}.$ (61)
+From (42c), we have
+$\frac{1}{n^{2}}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L\|{{x_{i,t}}-{x_{j,t}}}\|}]}}}\leq 2R(\mathbb{X})L{\vartheta_{1}}+\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{2R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}\|{\varepsilon_{i,t}^{z}}\|}]}}.$ (62)
+From (54), we have
+$\frac{{2{G_{1}}}}{n}\sum\limits_{i=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{\bar{z}}_{i,t}}-{{\bar{z}}_{t}}}\|}]}}\leq{G_{1}}{\vartheta_{1}}+\frac{{{G_{1}}{{\tilde{\vartheta}}_{2}}}}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{\|{\varepsilon_{i,t}^{z}}\|}]}}.$ (63)
+From (41), we have
+$\;\;\;\;\;\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{\varepsilon_{i,t}^{z}}\|}]}}$
+$=\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{{z_{i,t+1}}-{{\hat{z}}_{i,t+1}}}\|}]}}+\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}[{{{\tilde{\vartheta}}_{1}}\|{{z_{i,t+1}}-{x_{i,t}}}\|}]}}$
+$\leq\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}\sum\limits_{t=1}^{T}{{s_{t+1}}}+\frac{1}{n}\sum\limits_{t=1}^{T}{\sum\limits_{i=1}^{n}{{\mathbf{E}_{\mathcal{C}}}\Big{[}{\tilde{\vartheta}_{1}^{2}{\alpha_{t}}+\frac{{{{\|{{z_{i,t+1}}-{x_{i,t}}}\|}^{2}}}}{{4{\alpha_{t}}}}}\Big{]}}}.$ (64)
 Combining (61)–(62) and (42a), and noting that $\{{{s_{t}}}\}$ is nonincreasing, we know that (58a) holds.
@@
 $\geq\frac{1}{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{{[{{g_{i,t}}({{x_{j,t}}})}]}_{+}}}\|}^{2}}}]-G_{2}^{2}{\mathbf{E}_{\mathcal{C}}}[{{{\|{{x_{i,t}}-{x_{j,t}}}\|}^{2}}},$ (65)
-where the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
+where the second inequality holds due to the nonexpansive property of the projection $\cdot$$_{+}$}; and the last inequality holds due to (5).
 From ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$, we have
@@
 $\leq\frac{1}{n}\sum\limits_{i=1}^{n}{\sum\limits_{j=1}^{n}{\sum\limits_{t=1}^{T}{{\mathbf{E}_{\mathcal{C}}}[{\|{{{[{{g_{i,t}}({{x_{i,t}}})}]}_{+}}}\|+{G_{2}}\|{{x_{i,t}}-{x_{j,t}}}\|}]}}},$ (71)
-where the first inequality holds due to ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$; the second inequality holds due to the nonexpansive property of the projection ${[\cdot]_{+}}$; and the last inequality holds due to (5).
+where the first inequality holds due to ${g_{t}}(x)={\rm{col}}({{g_{1,t}}(x),\cdot\cdot\cdot,{g_{n,t}}(x)})$; the second inequality holds due to the nonexpansive property of the projection $\cdot$$_{+}$}; and the last inequality holds due to (5).
 From (42c), we have
@@
 (i) From (13), for any $T\in\mathbb{N}_{+}$, we have
-[…]
+$\sum\limits_{t=1}^{T}{{\alpha_{t}}}={\alpha_{0}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}}={\alpha_{0}}\Big{(}\sum\limits_{t=2}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}}+1\Big{)}\leq{\alpha_{0}}\Big{(}\int_{1}^{T}{\frac{1}{{{t^{{\theta_{1}}}}}}dt}+1\Big{)}\leq\frac{{{{\alpha_{0}}T^{1-{\theta_{1}}}}}}{{1-{\theta_{1}}}}.$ (73)
 Similar to the way to get (73), from (13) with ${\theta_{2}}\in({{\theta_{1}},1})$, for any $T\in\mathbb{N}_{+}$, we have
@@
 Combining (58a) and (73)–(76), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}}}{{1-{\theta_{2}}}}{T^{1-{\theta_{2}}}}$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\frac{{\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}}}{{1-{\theta_{2}}}}{T^{1-{\theta_{2}}}}$
 $\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}{T^{1+{\theta_{1}}-{\theta_{2}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (77)
 From (13) with ${\theta_{2}}=1$, for any $T\in\mathbb{N}_{+}$, we have
-[…]
+$\sum\limits_{t=1}^{T}{{s_{t}}}={s_{0}}\sum\limits_{t=1}^{T}{\frac{1}{t}}\leq{s_{0}}\Big{(}{\int_{1}^{T}{\frac{1}{t}dt}+1}\Big{)}\leq{s_{0}}({\log(T)+1})\leq 2{s_{0}}\log(T),{\rm{if}}\;T\geq 3$ (78)
 $\sum\limits_{t=1}^{T}{\frac{{{s_{t}}}}{{{\alpha_{t}}}}}=\frac{{{s_{0}}}}{{{\alpha_{0}}}}\sum\limits_{t=1}^{T}{\frac{1}{{{t^{1-{\theta_{1}}}}}}}\leq\frac{{{{s_{0}}T^{{\theta_{1}}}}}}{{{\theta_{1}}}{\alpha_{0}}}.$ (79)
 Combining (58a), (73), (76) and (78)–(79), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+4n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\log(T)$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}\log(T)$
 $\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{{\theta_{1}}}{\alpha_{0}}}{T^{{\theta_{1}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (80)
@@
 Combining (58a), (73), (76) and (81)–(82), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{1}}{s_{0}}$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{1}}{s_{0}}$
 $\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{({1+{\theta_{1}}-{\theta_{2}}}){\alpha_{0}}}{T^{1+{\theta_{1}}-{\theta_{2}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (83)
@@
 Combining (58a), (73), (76) and (84)–(85), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{2}}{s_{0}}$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{2}}{s_{0}}$
 $\;\;+\frac{{4{\tilde{p}}\sqrt{C}R(\mathbb{X}){s_{0}}}}{{{\alpha_{0}}}}\log(T)+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (86)
@@
 Combining (58a), (73), (76) and (87)–(88), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+2n{\tilde{p}}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{\rm Z_{3}}{s_{0}}$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{{\vartheta_{2}}{\alpha_{0}}}}{{1-{\theta_{1}}}}{T^{1-{\theta_{1}}}}+\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{\rm Z_{3}}{s_{0}}$
 $\;\;+\frac{{2{\tilde{p}}\sqrt{C}R(\mathbb{X}){\rm Z_{3}}{s_{0}}}}{{{\alpha_{0}}}}+\frac{{2R{{(\mathbb{X})}^{2}}}}{{{\alpha_{0}}}}{T^{{\theta_{1}}}}.$ (89)
@@
 Combining (58a), (76), (105)–(108), from the arbitrariness of ${y}\in{\mathcal{X}_{T}}$, we have
-$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq 2nR(\mathbb{X})L{\vartheta_{1}}+\frac{{2{\vartheta_{2}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{2n\tilde{p}\sqrt{C}R(\mathbb{X})L{{\tilde{\vartheta}}_{2}}{s_{0}}\mu}}{{1-\mu}}$
+$\mathbf{E}_{\mathcal{C}}[{{\rm{Net}\mbox{-}\rm{Reg}}(T)}]$ $\leq{\hat{\vartheta}}+\frac{{2{\vartheta_{2}}{\alpha_{0}}\mu}}{{1-\mu}}\sqrt{T}+\frac{{\tilde{p}\sqrt{C}{{\tilde{\vartheta}}_{1}}{s_{0}}\mu}}{{1-\mu}}$
 $\;\;+\frac{{2\tilde{p}\sqrt{C}R(\mathbb{X}){s_{0}}\sqrt{\mu}}}{{{\alpha_{0}}({1-\sqrt{\mu}})}}+\frac{{2R{{(\mathbb{X})}^{2}}({1-\mu})}}{{{\alpha_{0}}\mu}}.$ (109)